### v1

# OmniFusion: Simultaneous Multilingual Multimodal Translations via Modular Fusion

Abstract
There has been significant progress in open-source text-only translation large language models (LLMs) with better language coverage and quality. However, these models can be only used in cascaded pipelines for speech translation (ST), performing automatic speech recognition first followed by translation. This introduces additional latency, which is particularly critical in simultaneous ST (SimulST), and prevents the model from exploiting multimodal context, such as images, which can aid disambiguation. Pretrained multimodal foundation models (MMFMs) already possess strong perception and reasoning capabilities across multiple modalities, but generally lack the multilingual coverage and specialized translation performance of dedicated translation LLMs. To build an effective multimodal translation system, we propose an end-to-end approach that fuses MMFMs with translation LLMs. We introduce a novel fusion strategy that connects hidden states from multiple layers of a pretrained MMFM to a translation LLM, enabling joint end-to-end training. The resulting model, OmniFusion, built on Omni 2.5-7B as the MMFM and SeedX PPO-7B as the translation LLM, can perform speech-to-text, speech-and-image-to-text, and text-and-image-to-text translation. Experiments demonstrate that OmniFusion effectively leverages both audio and visual inputs, achieves a 1-second latency reduction in SimulST compared to cascaded pipelines and also improves the overall translation quality [^1] Code is available at https://github.com/saikoneru/OmniFusion.
OmniFusion: Simultaneous Multilingual Multimodal Translations via Modular Fusion
Sai Koneru1, Matthias Huck2, and Jan Niehues1 1 Karlsruhe Institute of Technology 2 SAP SE, Dietmar-Hopp-Allee 16, 69190 Walldorf, Germany {sai.koneru, jan.niehues}@kit.edu {matthias.huck}@sap.com
1 Introduction
In recent years, research in natural language processing (NLP) has increasingly shifted from unimodal foundation models (Touvron et al., 2023; Grattafiori et al., 2024) to multimodal foundation models (MMFMs) (Abdin et al., 2024; Goel et al., 2025; Xu et al., 2025), as many tasks benefit from contextual information derived from multiple sources. Although unimodal systems can perform strongly in certain settings, further progress also requires models to be able to take advantage of additional contextual cues without degrading quality.
One representative task that can benefit from such multimodal integration is speech translation (ST), where the goal is to translate spoken input in a source language into a target language. Over the years, advances in data scale, modeling techniques, and architectural design have led to substantial improvements in ST quality across many language pairs and domains (Ahmad et al., 2024; Agostinelli et al., 2025). However, the use of multimodal context—such as visual information from images or presentation slides accompanying speech—remains relatively underexplored (Liu and Niehues, 2024; Gaido et al., 2024a; Sinhamahapatra and Niehues, 2025).
Figure 1: Architecture of OmniFusion with supported tasks and inference modes.
Although MMFMs are capable of processing multiple modalities jointly, their multilingual translation capabilities are generally weaker than those of specialized translation large language models (LLMs) (Alves et al., 2024; Cheng et al., 2025). As a result, directly using MMFMs for ST often yields inferior translation quality compared to approaches that rely on strong text-based translation components (Appendix: Table 4). They also do not cover several languages compared to specialized translation LLMs. Consequently, a practical solution is to adopt a cascaded pipeline, where automatic speech recognition (ASR) is first performed, followed by machine translation (MT). This modular setup allows each component to be optimized independently and integrated during inference (Ahmad et al., 2024; Agostinelli et al., 2025; Koneru et al., 2025). However, since the translation model operates solely on text, it cannot exploit contextual information from the original speech or visual input. Moreover, the cascaded design introduces additional generation steps, leading to increased latency compared to end-to-end (E2E) models—an issue that is particularly critical for simultaneous ST (SimulST) applications (Agostinelli et al., 2025; Papi et al., 2025a).
Therefore, we aim to develop an E2E MMFM that excels in multimodality, multilinguality and efficient for SimulST. This can be achieved by extending the capabilities of a translation LLM to handle additional modalities. Few prior works have explored adding audio or image inputs to multilingual LLMs (Ambilduke et al., 2025; Viveiros et al., 2025). However, these approaches typically rely on training from scratch or only using pretrained audio/vision encoders and do not leverage the higher-level representational capabilities of existing MMFMs. Furthermore, training is ineffective as the base LLM needs to learn the modality mapping from scratch.
In this work, we propose leveraging pretrained MMFMs to introduce multimodal capabilities into a multilingual LLM with a E2E architecture. Rather than transferring only vision or audio encoders, we extract and utilize hidden-state representations from multiple layers of the MMFM to capture both the perceptual and reasoning components of multimodal understanding. Specifically, we adapt the translation LLM by incorporating hidden states from the first, middle, and last layers of the MMFM, thereby enabling the transfer of both low-level perception and high-level semantic understanding.
To further strengthen cross-modal alignment, we introduce Optical Character Recognition (OCR) as a bridge task for the image modality, analogous to how ASR serves as an alignment mechanism for audio. This strategy allows the model to effectively leverage pretrained multimodal representations while receiving direct cross-modal supervision.
We use Qwen Omni 2.5 7B (Xu et al., 2025) as the MMFM and SeedX 7B (Cheng et al., 2025) as the translation LLM to validate our approach on SimulST and multimodal MT tasks. We refer to the resulting fused model as OmniFusion.
Our main contributions are summarized as follows:
  • We propose a gated fusion approach that extracts multimodal embeddings from MMFMs, allowing translation LLMs to fully utilize the pretrained perceptual and interpretive capabilities of MMFMs.
  • We show the benefit of using slides for context and achieve an improvement in latency of approximately $1$ second in the SimulST scenario compared to cascaded fine-tuned models while having better quality. (Figure 2).
  • We observe a reduction in the major and critical errors for offline ST through self-cascading and effective use of the image context (Table 2).
  • We achieve state-of-the-art performance in the CoMMuTE benchmark for image-text translation highlighting the ability of the model to exploit images (Table 3).
2 Approach
We propose to fuse a MMFM with a task-specific translation LLM to leverage their complementary strengths and enable E2E training for improved latency. Furthermore, we aim to incorporate not only the perception capabilities of the MMFM but also its multimodal reasoning and understanding.
First, we describe the overall architecture and training process of our approach, considering a general input and output sequence independent of the specific task. We formally define the information flow, training loss, and the fusion of MMFM layers. Next, we detail the training tasks used to align multimodal representations from both image and audio modalities to the translation LLM. Finally, we discuss the training prompts that enable a self-cascading inference mode, allowing the model to enhance translation quality at the cost of increased latency. An overview of the architecture and supported tasks is in Figure 1.
2.1 Architecture
Let us denote the specialized translation LLM as $\mathcal{M}_{\text{NMT}}$ and the MMFM as $\mathcal{M}_{\text{MMFM}}$, where both use a language model as their backbone. The $\mathcal{M}_{\text{NMT}}$ is trained for translation via supervised fine-tuning with the following loss function:
$\mathcal{L}_{\text{Trans}}$ $=-\frac{1}{T_{y}}\sum_{t=1}^{T_{y}}\log P_{\theta}\big(y_{t}\mid y_{<t},\,X\big),$ (1)
$X$ $=(x_{1},\dots,x_{L_{x}}),\quad y=(y_{1},\dots,y_{T_{y}})$
where $X=(x_{1},\dots,x_{L_{x}})$ is the sequence of source tokens of length $L_{x}$, and $y=(y_{1},\dots,y_{T_{y}})$ is the sequence of target tokens of length $T_{y}$.
For $\mathcal{M}_{\text{MMFM}}$, the image and audio modalities are first processed by modality-specific encoders and then projected to match the embedding dimension of $\mathcal{M}_{\text{MMFM}}$, denoted as $D_{\text{MMFM}}$. This projection allows the encoder outputs to be concatenated as a sequence of tokens, enabling the MMFM to operate similarly to a standard LLM, such as $\mathcal{M}_{\text{NMT}}$.
To fuse both the perception and higher-level understanding of the MMFM, we hypothesize that lower layers predominantly capture perceptual features, while higher layers encode more abstract reasoning as computation progresses. Accordingly, we extract the first, middle, and last hidden layers from the MMFM, denoted as:
$h^{(1)}_{t}$ $=\text{MMFM}_{\text{layer 1}}(x_{t}),$ (2)
$h^{(\text{mid})}_{t}$ $=\text{MMFM}_{\text{layer mid}}(x_{t}),$
$h^{(\text{last})}_{t}$ $=\text{MMFM}_{\text{layer last}}(x_{t}),$
$h^{(1)}_{t},\,h^{(\text{mid})}_{t}$, $\,h^{(\text{last})}_{t}\in\mathbb{R}^{L_{m}\times D_{\text{MMFM}}},t=1,\dots,L_{m}$
where $L_{m}$ is the length of the MMFM hidden states.
Directly concatenating these multi-dimensional representations into $\mathcal{M}_{\text{NMT}}$ poses three challenges. First, the sequence length triples, significantly increasing computational and memory requirements. Second, the relevance of each layer may vary depending on the task, necessitating selective weighting. Third, the embedding dimension of $\mathcal{M}_{\text{MMFM}}$ may not match that of $\mathcal{M}_{\text{NMT}}$.
To overcome these challenges, we introduce a gated fusion layer that determines the contribution of each hidden state along the sequence length. Each hidden state is multiplied by its corresponding gate values, summed across layers, and then projected through an MLP to align the embedding dimension with that of the translation model:
$\begin{aligned} W_{\text{gate}}&\in\mathbb{R}^{3\times 3D_{\text{MMFM}}},\\ [g_{1},g_{\text{mid}},g_{\text{last}}]_{t}&=\text{softmax}\big(W_{\text{gate}}\cdot h^{(1)}_{t}\oplus h^{(\text{mid})}_{t}\oplus h^{(\text{last})}_{t}\big),\\ m^{\prime}_{t}&=g_{1}\odot h^{(1)}_{t}+g_{\text{mid}}\odot h^{(\text{mid})}_{t}+g_{\text{last}}\odot h^{(\text{last})}_{t},\\ m_{t}&=\text{MLP}(m^{\prime}_{t})\in\mathbb{R}^{L_{m}\times D_{\text{Trans}}},\quad t=1,\dots,L_{x}\end{aligned}$
where $D_{\text{Trans}}$ is the embedding dimension of $\mathcal{M}_{\text{NMT}}$ and $m_{t}$ is the fused multimodal representation at time step $t$.
The fused multimodal embeddings $m_{t}$ can now be treated like additional token embeddings and concatenated with the $\mathcal{M}_{NMT}$ token embeddings:
$\tilde{x}_{t}$ $=m_{t}\oplus x_{t},$ (4)
$\tilde{X}$ $=(\tilde{x}_{1},\dots,\tilde{x}_{L_{m}+L_{x}})$
Finally, the multimodal translation loss is computed by conditioning on the concatenated sequence $\tilde{X}$ while predicting the target tokens:
[…]
Where $\tilde{X}$ is the multimodal input to the translation model.
2.2 Training
We have shown how MMFM internal representations can be extracted and integrated into our downstream model. However, leveraging these representations effectively also requires a training process that explicitly aligns multimodal features so that the LLM can exploit them. The exact alignment procedure depends on the downstream tasks we expect the fused model to perform after fine-tuning.
In this work, our goal is to obtain a translation model that can utilize both audio and image context. Therefore, we must make the translation LLM context-aware by training to exploit these representations. We consider $3$ translation tasks in this work and are defined as follows:
  • Speech Translation (ST): Generate a translation in the target language given only audio in the source language.
  • Speech-Image Translation (SIT): Generate a translation in the target language given source-language audio together with an image that provides additional context (e.g., a slide corresponding to the spoken segment).
  • Text-Image Translation (TIT): Generate a translation in the target language given the source text and an aligned image, such as in caption translation scenarios.
While we can train the model to perform translation directly, decomposing the task into intermediate steps can be beneficial. For example, translating English audio to German can be approached by first predicting ASR outputs and then producing the translation by continuing generation conditioned on the ASR hypothesis. This decomposition not only simplifies learning but also provides explicit alignment signals that help map audio representations into the LLM’s semantic space.
When vision input is available—particularly for slide-based scenarios—we additionally introduce OCR as a bridge task before generating the final translation. In this case, the model must predict both OCR text and the final translation using the same audio–visual representation. This encourages alignment between the modalities and explicitly informs the model that the visual context is meaningful and should be utilized.
During training, we stochastically decide—based on a specified probability for each data source—whether an example should be trained in self-cascading mode and along which modality the cascade should occur. Furthermore, we provide modality-specific prompt to the MMFM for handling different modality combination scenarios. Pseudocode for building the prompt for each task during training is shown in Algorithm˜1.
We also freeze the MMFM to preserve its pretrained abilities while fine-tuning the translation LLM with LoRA (Hu et al., 2021). We use LoRA for computational benefits and not go Out-Of-Memory during fine-tuning.
3 Experimental Setup
3.1 Training Data
We aim to train the fused model to perform translation tasks across both speech and image modalities. An overview of the data sources is provided in Table˜1.
For ST, we use subsets of Europarl-ST (Iranzo-Sánchez et al., 2020) and Covost-2 (Wang et al., 2020), sampling across multiple target languages. We limit training data to these subsets due to computational constraints and to demonstrate that full datasets are not strictly necessary.
To enable the model to exploit images when translating audio, we use the M3AV corpus (Chen et al., 2024), which contains English scientific talks paired with corresponding slides. The slides also include Paddle-OCR (Cui et al., 2025) predictions, allowing the model to be trained either to predict translations directly or via an intermediate OCR step for better alignment.
Since M3AV has limited language coverage, we augment the data by synthetically translating ASR transcripts into additional target languages (M3AV-Aug). This enables the model to handle ST with slides and supports the self-cascade mode, performing ASR first and then translation.
Finally, for TIT, we employ the Multi30k dataset (Elliott et al., 2016), which contains images paired with captions in multiple languages. This task encourages the model to go beyond OCR and to leverage visual content, including objects and scene understanding, for translation.
3.2 Models
Several MMFMs have been developed rapidly in recent years, showing significant improvements in both quality and efficiency (Abdin et al., 2024; Team et al., 2025). In our scenario, we require a model capable of processing both audio and image modalities, while also supporting cross-modal abilities. Accordingly, we select Qwen Omni 2.5B [^2] Qwen/Qwen2 Audio.5-Omni-7B (Xu et al., 2025), due to its strong multimodal capabilities and the potential to incorporate video in future experiments.
For the translation LLM, we choose the recently proposed SeedX PPO 7B [^3] ByteDance-Seed/Seed-X-PPO-7B (Cheng et al., 2025), which provides broad language coverage and strong translation performance compared to closed-source alternatives.
We use the training data mentioned above and denote the fine-tuned model as OmniFusion for the rest of the paper. While we report results using these specific models, it is important to note that our approach is not limited to these base LLMs. The proposed fusion strategy does not require massive pre-training or fine-tuning resources and can be adapted in more modest computing environments, such as university clusters. The training hyper-parameters are specified in Table˜10. For validation, we used Covost-2 validation data and select the checkpoint with the least loss in 20k steps.
Dataset # Examples Language Modality Task
M3AV 122k en Speech, Image ASR, OCR
M3AV - Aug 122k
zh, fr, de, it,
ja, ko, pt, ru,
es, vi, ar, cs,
hr, da, nl, fi,
hu, id, ms, nb,
no, pl, ro, tr
Speech, Image ST, OCR
Europarl 260k*
de, es, fr, it, nl,
pl, pt, ro
Speech ST
Covost2 de, ja, tr, zh Speech ST
Coco-Captions 87k cs, de, fr Image TIT
Table 1: Overview of data sources collected for E2E training. * We sample together 260k from the concatenated data uniformly across the language pairs. All the speech and text data are only available with English as the source language. Language code mapping is presented in Table˜10.
3.3 Evaluation: Data & Metrics
For evaluating the performance of multimodal Simul-ST systems, we require aligned audio-visual translation data. Additionally, we avoid using M3AV for testing, as it overlaps with our training data, monolingual and we aim to measure the model’s robustness on unseen domains.
Consequently, we use the MCIF test set (Papi et al., 2025b), which contains video and audio recordings of ACL talks in three target languages from English. We report latency with Average Lagging (Ma et al., 2020) and translation quality results on this dataset.
For TIT, we focus on test cases where the image is crucial for disambiguating the source sentence. For this, we use the CoMMuTE dataset (Futeral et al., 2023, 2025), which provides a small but multilingual targeted TIT test set.
We measure translation quality using XCOMET-XL (Guerreiro et al., 2024) for ST scenarios, as it offers high-quality evaluation and allows analysis of minor, major, and critical errors. For TIT, we use the COMET [^4] Unbabel/wmt22-comet-da model (Rei et al., 2022) directly, without providing error spans, to evaluate aligned references.
4 Results
Our primary motivation for the proposed approach is to build an E2E ST system that is context-aware, efficient for SimulST and enabled to exploit additional image context. To validate this, we first evaluate the system in simultaneous scenarios by comparing latency and quality. Next, we report quality in offline scenarios and conduct an error analysis to contrast the E2E system with traditional cascaded systems fine-tuned on the same data. We then assess performance in TIT scenarios, benchmarking against other off-the-shelf multimodal translation models. Finally, we perform an ablation study to analyze the contributions of individual layers in the MMFM and present the key findings.
We also validate the proposed approach for a different model combination (Qwen-Audio (Chu et al., 2023) and Tower-Instruct (Alves et al., 2024)), analyze training with different layers and post-editing in our preliminary experiments and report the results in the Appendixes Table˜4 and table˜5.
(a) en $\rightarrow$ de
(b) en $\rightarrow$ it
(c) en $\rightarrow$ de
(d) en $\rightarrow$ it
[…]
4.1 Simultaneous Speech Translation
A SimulST system must generate translations incrementally as the source audio arrives, rather than waiting for the entire utterance to finish. The decision of whether to commit a token or read more audio is referred to as the policy. Since our E2E system does not natively support simultaneous translation, we adopt the widely used Local Agreement policy (Liu et al., 2020), which commits the longest common prefix observed across consecutive predictions. Concretely, at each step we process the audio in fixed chunk sizes (a hyperparameter) and irreversibly emit the overlapping prefix between the current and previous hypotheses. After committing, subsequent tokens are force-decoded conditioned on the partial translation committed so far.
To fairly evaluate the proposed approach, we compare our E2E model OmniFusion against a strong cascaded counterpart built from the same pretrained components. We therefore fine-tune this cascaded baseline on the same data using LoRA adapters (Omni + SeedX LoRA FT). The effect of fine-tuning is discussed in more detail in Section˜4.2. We experiment with chunk sizes of $[1,1.5,2,2.5,3]$ seconds, run inference both with and without images, and report computation-aware and computation-unaware AL together with XCOMET-XL scores for the $en\rightarrow de$ and $en\rightarrow it$ language pairs in Figure˜2.
The results reveal several key insights. First, as expected, the offline ST systems (with images) achieve the highest quality in both language pairs due to access to full context. Next, OmniFusion is consistently around 1 second faster than its cascaded counterpart across all chunk sizes while preserving comparable quality in the audio-only setting. When images are provided, OmniFusion delivers both the lowest latency and the highest quality, demonstrating that visual context is substantially more beneficial in the E2E model than in the cascaded baseline allowing it to commit high quality stable outputs earlier. Finally, we also see that in computation-aware settings, adding image causes delay due to processing and causes overhead than only-audio creating a trade-off between quality and latency.
We attribute the advantage of OmniFusion to two factors: (1) our fine-tuning strategy, which includes an OCR objective that tightly aligns image and audio representations, and (2) the inherent efficiency of the E2E architecture, which requires fewer decoding steps overall. In contrast, the cascaded pipeline with Local Agreement must wait for stable ASR hypotheses before translation can even begin, making it slower than the E2E approach.
In conclusion, OmniFusion achieves better latency, leverages image context effectively, and has higher quality in SimulST scenarios, making it particularly well-suited for real-world applications such as live presentations and meetings where visual information is readily available.
4.2 Offline Speech Translation
Category Config Image XCOMET-XL Minor Major Critical Total
Cascade Systems (ASR → MT)
Omni + SeedX
No FT 85.88 1212.3 773.0 74.3 2059.7
No FT 85.94 1191.3 759.3 74.7 2025.3
LoRA FT 86.42 1239.0 751.3 75.0 2065.3
LoRA FT 86.59 1202.3 766.3 55.7 2024.3
End-to-End (Direct ST)
OmniFusion
Mid Fusion 83.57 1114.0 1188.3 72.3 2374.7
Mid Fusion 85.98 1256.3 749.0 61.7 2067.0
Gated Fusion 83.98 1119.3 1188.0 62.7 2370.0
Gated Fusion 86.24 1255.7 736.0 57.0 2048.7
Gated Fusion + Self-Casade 86.57 1231.0 719.0 55.3 2005.3
[…]
Although SimulST is the primary focus of this work, we also evaluate the systems in the offline scenario, where latency is not a concern and the goal is to maximize translation quality. In this setting, we can leverage the self-cascade inference mode described in Section˜2.2 to further boost performance.
To measure the contributions of fine-tuning and visual context, we first evaluate the cascaded baseline in four configurations: base vs. fine-tuned (LoRA FT) models, each tested with and without images. For the E2E OmniFusion model, we additionally train a variant that uses only the middle-layer fusion without gating to assess the need for the gating module. We then compare all E2E configurations (mid-only fusion and full gated fusion) both with and without images, and with/without self-cascade inference. Results are averaged across three target languages $en\rightarrow\{de,it,zh\}$ and reported in Table˜2 using XCOMET-XL scores. We further include the detailed XCOMET-XL error categories, hypothesizing that E2E models may sacrifice minor fluency for substantially fewer major and critical errors.
The results yield several insights. In the cascaded pipeline, LoRA fine-tuning yields consistent gains (e.g., 85.94 $\rightarrow$ 86.59 with images), whereas adding images provides only marginal benefit. In contrast, the E2E OmniFusion model benefits substantially from both gated fusion (outperforming mid-only fusion) and image context showing the need for gating module and additional contextual information. Although the direct E2E translation score remains slightly below the best cascaded system, enabling self-cascade inference closes this gap almost entirely (cascaded LoRA FT: 86.59 vs. E2E gated fusion + self-cascade: 86.57). More importantly, error analysis reveals that our best E2E configuration produces the fewest critical and major errors (e.g., 719 vs. 751.3 for the strongest cascaded baseline). These findings demonstrate that OmniFusion, when equipped with image context and self-cascade inference, matches fine-tuned cascaded system in offline ST quality while offering significantly fewer severe translation errors
4.3 Text-Image Translation
Model Config en $\rightarrow$ ar en $\rightarrow$ cs en $\rightarrow$ de en $\rightarrow$ fr en $\rightarrow$ ru en $\rightarrow$ zh
Omni 2.5B - 78.56 81.03 81.94 79.99 47.27 81.74
SeedX 7B - 80.39 83.56 82.38 80.31 82.13 83.25
ZeroMMT 3.3B
(Futeral et al., 2025)
- 80.64 86.69 83.54 83.40 83.87 77.85
TowerVision 9B
(Viveiros et al., 2025)
- 70.85 86.62 84.97 84.29 85.81 85.24
OmniFusion 14B Mid Fusion 80.98 87.49 85.14 84.97 85.69 85.15
OmniFusion 14B Gated Fusion 80.81 87.58 84.88 84.28 84.62 84.34
Table 3: COMET scores ($\uparrow$) on the CoMMuTE dataset. Statistically significant top-performing results per language pair are highlighted in bold. ‘Config‘ distinguishes the two OmniFusion variants (mid-layer fusion vs. gated fusion).
Apart from ST, translating captions also relies on image context, where cascaded systems fail due to their inability for disambiguation. To demonstrate this, we evaluate several models on the CoMMuTE test set across multiple language pairs and report COMET scores in Table˜3.
First, the audio-only Omni 2.5B exhibits the weakest multilingual performance, which reinforces our motivation for integrating it with a strong translation backbone. SeedX 7B obtains reasonable scores yet consistently trails all vision-capable models, confirming its limitation in the absence of image context. Among vision-aware systems, both Mid Fusion and Gated Fusion variants of our OmniFusion 14B achieve state-of-the-art or near-SOTA COMET scores, matching or surpassing vision-only models such as ZeroMMT 3.3B (Futeral et al., 2025) and TowerVision 9B (Viveiros et al., 2025) on most language directions.
Interestingly, and in contrast to ST, the Mid Fusion variant slightly outperforms Gated Fusion on CoMMuTE. This suggests that, when the image serves primarily as supplementary context (as in caption translation) rather than the main information source (as in audio for spoken ST), using visual features at the middle layer alone may be sufficient, and the additional flexibility of gating offers no further advantage.
Importantly, unlike TowerVision which requires massive multimodal pretraining from scratch, OmniFusion can be built on top of existing strong pretrained components. This makes our approach far more data- and training-efficient without compromising quality, by effectively exploiting the rich image–text alignment knowledge already present in the MMFM pretrained weights. These results underline the strength of our E2E design on visually grounded translation tasks where image context is essential.
4.4 Contribution of Layers
Figure 3: Analysis of the layer contribution from the MMFM in the gating module across modalities.
We presented a gating module in Section˜2.1 for fusing information across layers of the MMFM. To better understand its behavior, we conduct an ablation study by examining the gate values assigned to each MMFM layer during fusion. Specifically, we compute the average gate value across the three layers (first, middle, last), noting that the gate weights for each timestep sum to 1.
We run inference on 100 examples from the MCIF dataset—both speech-only and speech+image conditions—and on the CoMMuTE dataset for the image-only modality. The resulting average gate values are shown in Figure˜3.
We observe that the first and middle layers receive consistently higher contributions, while the last layer contributes only marginally. This suggests that the majority of transferable multimodal processing occurs in the earlier half of the MMFM. Although the weak contribution of the last layer is not entirely clear, we hypothesize that these representations may be more language-specific, whereas the middle layers capture more language-agnostic features (Liu and Niehues, 2025).
5 Related Work
Multimodal MT: Translating captions has been widely explored in the MT field over the last decade (Specia et al., 2016; Elliott et al., 2017; Barrault et al., 2018). Several works focus on enabling NMT models to process images, either by training models from scratch (Vijayan et al., 2024) or by augmenting large pretrained models such as NLLB (Costa-Jussà et al., 2022; Futeral et al., 2025). Most recently, and in parallel to our work, Viveiros et al. (2025) extend Tower with image and video modalities through a computationally expensive multi-stage training pipeline.
While these efforts focus primarily on vision, an other line of direction is adding audio to translation LLMs (Ambilduke et al., 2025). A smaller line of work tackles the combination of vision and audio, such as Sinhamahapatra and Niehues (2025), which exploits slides and speech for ASR. Building on this trend, there is also increasing interest in translating audio together with video (Anwar et al., 2023; Choi et al., 2024). In contrast to these works, our approach focuses on jointly using images and audio specifically for ST, while remaining efficient in both training and inference.
Simultaneous ST: Many real-world scenarios require real-time ST without waiting for the full source utterance. However, most research focuses on training models in offline settings and only later adapting them to SimulST. Such adaptation typically relies on predefined or learned translation policies, such as wait-k (Ma et al., 2019; Elbayad et al., 2020), which assumes a fixed token delay before committing to output. Other approaches include Local Agreement (Liu et al., 2020), which selects the most stable prefix across consecutive read steps or attention-based heuristics (Papi et al., 2023). More recently, several works explore leveraging LLMs for SimulST policy learning (Koshkin et al., 2024b, a; Guo et al., 2025; Fu et al., 2025). In our work, we use Local Agreement to validate the effectiveness of our approach and to compare against cascaded baselines. However, our method is compatible in theory with other SimulST policies, and integrating more advanced or LLM-driven strategies is a promising direction for future work.
Modality Fusion: With the rapid development of multimodal foundation models, a key challenge is how to effectively integrate speech and image modalities into text-centric LLMs. Some approaches (Abdin et al., 2024; Xu et al., 2025; Team et al., 2025) process each modality with separate encoders and merge them later within the LLM backbone. In contrast, models such as (Ye et al., 2025) enable early interaction between vision and audio modalities, facilitating better cross-modal alignment. Beyond the timing of modality interaction, ongoing research also investigates the design of modality projectors (Verdini et al., 2024; Gaido et al., 2024b) and the use of continuous versus discrete tokens for non-text modalities (Zhan et al., 2024; Li et al., 2025), balancing the trade-offs between latency and translation quality.
6 Conclusion
We proposed a novel approach to fuse a MMFM with a translation LLM, and validated its effectiveness across three key scenarios. First, we showed that the fused model achieves faster inference while maintaining strong translation quality in SimulST settings, even when only partial source audio is available. Second, the model produces fewer major and critical translation errors, demonstrating more reliable linguistic accuracy. Third, we found that the model effectively leverages visual context for both speech and text translation, leading to consistent gains in multimodal settings. Finally, our layer-level analysis reveals that the first and middle MMFM layers contribute most to capturing semantic multimodal information, providing insight into where meaningful cross-modal representations emerge.
For future work, our fusion framework can naturally be extended to additional modalities. Since Omni already supports video processing, incorporating video embeddings would enable richer context for translation. Moreover, the fusion strategy is not limited to models for multimodal inputs but future extensions could target multimodal generation, allowing the system to produce not only text but also speech, images, or video as output.
Limitations
A primary limitation of our work is that OmniFusion is trained with fixed task-specific prompts across the three supported translation settings, and therefore does not exhibit general instruction-following capabilities. It remains unclear whether our fusion strategy can be extended to support more flexible, instruction-driven multimodal behavior. In addition, most of our experiments use English as the source language. Although we report zero-shot results in Table 4, we do not evaluate the OmniFusion model with vision inputs in multilingual zero-shot conditions.
Appendix A Appendix
Model Mode Hypothesis Comet-22 ($\uparrow$)
[…]
Baselines
Qwen2 Audio E2E _ 81.24 78.20 73.59 83.38 70.70 71.26
Qwen2 Audio FT E2E _ 84.82 85.57 80.22 84.17 67.69 77.09
Qwen2 Audio + Tower Cascade _ 85.85 85.97 82.94 85.36 69.26 84.78
Tower APE Cascade APE Qwen2 Audio + Tower 85.96 85.77 82.59 85.21 47.43 70.67
Off-the-shelf ST models
Tower-Spire E2E _ 82.77 81.63 _ _ _ _
Tower-Spire Self-Cascade _ 85.17 85.73 _ _ _ _
Seamless-M4T v2 E2E _ 85.63 80.05 70.22 77.78 70.58 78.91
E2E models
AudioTower-Zero E2E _ 85.13 83.97 76.62 83.98 68.71 81.64
AudioTower-Mid E2E _ 86.02 85.03 80.62 84.86 74.47 83.26
AudioTower-Last E2E _ 82.97 81.03 54.12 49.57 50.19 52.31
E2E + Speech-Aware APE model
AudioTower-Mid Cascade APE Qwen2 Audio + Tower 86.34 85.99 83.42 85.96 78.26 84.59
AudioTower-Mid E2E Empty 85.77 84.79 80.39 85.27 74.67 82.71
Table 4: Offline ST results on FLEURS test set. Speech-Aware APE models use the hypothesis from a system and perform post-editing. Statistically significant top-performing results per language pair are highlighted in bold. Underlined langauges are not seen during the fusing process as source/target.
Model Mode Hypothesis Comet-22 ($\uparrow$)
[…]
Qwen2 Audio + Tower Cascade _ 85.85 85.97 82.94 85.36 69.26 84.78
Whisper + Tower Cascade _ 86.71 86.31 82.23 86.54 83.56 85.29
AudioTower-Mid Cascade APE Qwen2 Audio + Tower 86.34 85.99 83.42 85.96 78.26 84.59
AudioTower-Mid Cascade APE Whisper + Tower 87.09 86.34 82.81 86.60 84.38 84.95
Character-Error-Rate ($\downarrow$)
Whisper ASR _ 3.14 3.15 25.37 2.11 7.67 3.28
Table 5: Analysis of system combination by using Whisper to generate the hypothesis for post-editing on FLEURS test set. Statistically significant top-performing results per language pair are highlighted in bold. Underlined langauges are not seen during the fusing process as source/target.
A.1 Layer Analysis, Zero-shot ability and APE - AudioTower
In our preliminary experiments, we focus exclusively on the audio modality and investigate fusing Qwen-Audio 7B (Chu et al., 2023) with Tower-Instruct 7B v0.2 (Alves et al., 2024). We train the fused model for ST using the same setup described earlier, but without gating, and we independently fuse the first, middle, and last layers of Qwen-Audio to assess their individual contributions. We then compare this fused model against several baselines. First, we evaluate both the out-of-the-box and fine-tuned versions of Qwen-Audio (Qwen2 Audio FT), trained end-to-end on the same ST data. We also include a standard cascaded setup using off-the-shelf ASR and MT models. Additionally, we evaluate an Automatic Post-Editing (APE) configuration in which Tower corrects translations generated by Qwen-Audio.
To contextualize performance, we further compare against other off-the-shelf ST systems such as Tower-Spire (Ambilduke et al., 2025) and Seamless-M4T (Barrault et al., 2023), which provides insight into multilingual generalization. Beyond the main fused model, we also train an APE variant of our fusion architecture that post-edits the outputs of cascaded systems. Unlike traditional APE, this Speech-Aware APE allows the model to use the source audio when correcting errors, enabling more informed revisions. For both APE variants, we generate synthetic APE tuples by running Qwen-Audio on the training data and fine-tune with LoRA.
We report results on the FLEURS test set in Table 4, including evaluations on zero-shot language directions not present on either the source or target side during training. First, while fine-tuning improves Qwen2 Audio, it still falls short of the out-of-the-box cascaded setup across all languages, underscoring the strength of translation-centric LLMs. Tower-based APE is competitive for seen languages but struggles in harder zero-shot directions, such as frkr. Off-the-shelf end-to-end ST models lag significantly behind cascaded systems as well.
Our proposed fused model (AudioTower) with middle-layer fusion achieves the strongest overall performance and even surpasses the cascaded baseline in several languages. Notably, both the first and middle layers show strong transfer to unseen languages, whereas the last layer collapses entirely—supporting the hypothesis that upper layers encode highly language-specific features, making them less suitable for cross-lingual contextual embeddings.
Finally, the APE variant of the fused model delivers the best results overall by combining the strong hypotheses of the cascaded pipeline with the ability to exploit the raw audio when correcting errors. We also test this APE model with empty hypotheses and observe performance comparable to AudioTower-Mid, indicating its flexibility for multiple use cases.
Given the strong APE results, we additionally explore whether the model can correct outputs from different ASR systems. Using Whisper (Radford et al., 2023) for ASR, we evaluate the fused APE model on Whisper+Tower outputs and report results in Table˜5. When Whisper+Tower produce the strongest cascaded baseline, using them as hypotheses likewise yields the highest APE performance—highlighting that Speech-Aware APE can be beneficial not only for E2E setups but also for offline cascaded scenarios.
A.2 Prompt Building
System Prompt for all modalities You are Qwen, a virtual human developed by the Qwen Team, Alibaba Group, capable of perceiving auditory and visual inputs, as well as generating text and speech.
User Prompt (ST with/without Image) Transcribe the audio using the image for context, and perform OCR on the image for correct spelling of keywords and names.
User Prompt (TIT) Translate the following sentence into {tgt_lang} using the image for context. <IMG> refers to the image. Use the image to determine formality, gender, keywords, and other details important for disambiguation: {ex[src]} <IMG>
1:Example exex
2:// SeedX prompt (source text, language tag and reference)
[…]
5:
6:if aud and not img then $\triangleright$ Speech only
7:  if $\textsc{Random}<0.1$ then
8:   $\texttt{inp}\leftarrow\langleen\rangle$ $\triangleright$ English-Audio only in our dataset
9:   $\texttt{lbl}\leftarrow\textsc{ASR}(\texttt{src}){+}\langleeos$$\rangle{+}\texttt{suffix}{+}\texttt{tgt}{+}\langleeos\rangle$ $\triangleright$ ASR Self-cascade
10:  else
11:   $\texttt{inp}\leftarrow\texttt{suffix}$
12:   $\texttt{lbl}\leftarrow\texttt{tgt}{+}\langleeos\rangle$
13:  end if
14:else if img and not aud then $\triangleright$ Image only
15:  if $\textsc{Random}<0.05$ then
16:   Swap($\texttt{src},\texttt{tgt}$)
17:   $\texttt{suffix}\leftarrow\langleEnglish\rangle$
18:  end if
19:  $\texttt{src}_{m}\leftarrow\textsc{Mask}(\texttt{src},\texttt{prob\_words}=0.2,\texttt{mask\_chance}=0.2)$ $\triangleright$ Mask 20% source words for 20% of the time to pay attention to Image
[…]
21:  $\texttt{lbl}\leftarrow\texttt{tgt}{+}\langleeos\rangle$
22:else if img and aud then $\triangleright$ Audio + Image
23:  if $\textsc{Random}<0.8$ then
24:   if $\textsc{Random}<0.1$ then
25:     $\texttt{inp}\leftarrow\langleen\rangle$
26:     $\texttt{lbl}\leftarrow\textsc{ASR}(\texttt{src}){+}\langleeos$$\rangle{+}\texttt{suffix}{+}\texttt{tgt}{+}\langleeos\rangle$
27:   else
28:     $\texttt{inp}\leftarrow\texttt{suffix}$
29:     $\texttt{lbl}\leftarrow\texttt{tgt}{+}\langleeos\rangle$
30:   end if
31:  else
32:   $\texttt{inp}\leftarrow\langleOcr\rangle$ $\triangleright$ OCR Self-cascade
33:   $\texttt{lbl}\leftarrow\texttt{ocr}{+}\langleeos$$\rangle{+}\texttt{suffix}{+}\texttt{tgt}{+}\langleeos\rangle$
34:  end if
35:else
36:  error “no modalities”
37:end if
38:
39:// Omni prompt (Multimodal prompt builder)
40:if aud and img then
41:  $\texttt{user}\leftarrow\texttt{``Transcribe with OCR and image context:''}\;\|\;\texttt{img}\;\|\;\texttt{aud}$
42:else if aud then
43:  $\texttt{user}\leftarrow\texttt{``Transcribe audio:''}\;\|\;\texttt{aud}$
44:else
45:  $\texttt{user}\leftarrow\texttt{``Translate with image:''}\;\|\;\texttt{img}$
46:end if
[…]
Algorithm 1 Online Prompt building for examples in the training data to enable alignment and self-cascading inference mode.
Category Model Name Image XCOMET-XL Minor Major Critical Total
Cascade Systems (ASR → MT)
Omni + SeedX
No FT 89.66 2172 132 68 2372
No FT 89.49 2115 133 76 2324
LoRA FT 89.75 2209 128 67 2404
LoRA FT 90.18 2135 118 56 2309
End-to-End (Direct ST)
OmniFusion
Mid Fusion 88.44 2126 157 74 2357
Mid Fusion 90.03 2256 104 68 2428
Gated Fusion 88.09 2133 181 87 2401
Gated Fusion 89.90 2217 106 60 2383
Gated Fusion + Self-Cascade 90.05 2191 94 62 2347
[…]
Category Model Name Image XCOMET-XL Minor Major Critical Total
Cascade Systems (ASR → MT)
Omni + SeedX
No FT 86.03 850 932 51 1833
No FT 86.11 862 891 67 1820
LoRA FT 86.76 884 855 54 1793
LoRA FT 86.85 868 855 55 1778
End-to-End (Direct ST)
OmniFusion
Mid Fusion 82.80 754 1065 93 1912
Mid Fusion 85.75 899 903 64 1866
Gated Fusion 84.06 745 1085 56 1886
Gated Fusion 86.79 914 847 48 1809
Gated Fusion + Self-Cascade 87.20 879 838 44 1761
[…]
Category Model Name Image XCOMET-XL Minor Major Critical Total
Cascade Systems (ASR → MT)
Omni + SeedX
No FT 81.94 615 1255 104 1974
No FT 82.23 597 1254 81 1932
LoRA FT 82.74 624 1271 104 1999
LoRA FT 82.74 604 1326 56 1986
End-to-End (Direct ST)
OmniFusion
Mid Fusion 79.47 462 1343 50 1855
Mid Fusion 82.15 614 1240 53 1907
Gated Fusion 79.78 480 1298 45 1823
Gated Fusion 82.03 636 1255 63 1954
Gated Fusion + Self-Cascade 82.47 623 1225 60 1908
[…]
Language Code
Arabic ar
Chinese zh
Croatian hr
Czech cs
Danish da
Dutch nl
Finnish fi
French fr
German de
Hungarian hu
Indonesian id
Italian it
Japanese ja
Korean ko
Malay ms
Norwegian no
Norwegian Bokmål nb
Polish pl
Portuguese pt
Romanian ro
Russian ru
Spanish es
Turkish tr
Vietnamese vi
Table 9: Language codes used in paper following Cheng et al. (2025).
Hyperparameter Value
Gradient checkpointing True
Gradient checkpointing kwargs {’use_reentrant’: False}
Save steps 1000
Evaluation steps 1000
Save strategy steps
Evaluation strategy steps
Logging steps 5
Logging strategy steps
Learning rate 1e-4
Train batch size (per device) 4
Eval batch size (per device) 4
Gradient accumulation steps 2
Weight decay 0.01
Save total limit 3
Max steps 20000
bf16 True
Push to hub False
Metric for best model loss
Remove unused columns False
Label names ["labels"]
Dataloader num workers 1
DDP find unused parameters True
Table 10: Huggingface trainer hyperparameters used for the model on 4 Nvidia A100 48GB GPU’s with DDP training mode. Unspecified hyperparameters are set to default
Image Cascaded FT Output OmniFusion Output Reference
[Uncaptioned image] Hallo, willkommen zu unserer Präsentation von Deplain, einem neuen Korpus für die Identifizierung deutscher Texte auf Dokumenten- und Satzebene. Hallo, willkommen zu unserer Präsentation von DEPLAIN, einem neuen Korpus für die deutsche Texterkennung auf Dokumentebene und auf Satzebene. Herzlich willkommen zu unserer Präsentation von DEPLAIN, einem neuen Korpus für die deutsche Texterkennung auf Dokument- und Satzebene.
Table 11: Example illustrating where OmniFusion is better than cascaded variant. Example taken from MCIF segment 1.

### v2

@@
-OmniFusion: Simultaneous Multilingual Multimodal Translations via Modular Fusion
+OmniFusion: Simultaneous Multilingual and Multimodal Translations via Modular Fusion
 Abstract
-There has been significant progress in open-source text-only translation large language models (LLMs) with better language coverage and quality.
-However, these models can be only used in cascaded pipelines for speech translation (ST), performing automatic speech recognition first followed by translation.
-This introduces additional latency, which is particularly critical in simultaneous ST (SimulST), and prevents the model from exploiting multimodal context, such as images, which can aid disambiguation.
-Pretrained multimodal foundation models (MMFMs) already possess strong perception and reasoning capabilities across multiple modalities, but generally lack the multilingual coverage and specialized translation performance of dedicated translation LLMs.
-To build an effective multimodal translation system, we propose an end-to-end approach that fuses MMFMs with translation LLMs.
-We introduce a novel fusion strategy that connects hidden states from multiple layers of a pretrained MMFM to a translation LLM, enabling joint end-to-end training.
-The resulting model, OmniFusion, built on Omni 2.5-7B as the MMFM and SeedX PPO-7B as the translation LLM, can perform speech-to-text, speech-and-image-to-text, and text-and-image-to-text translation.
-Experiments demonstrate that OmniFusion effectively leverages both audio and visual inputs, achieves a 1-second latency reduction in SimulST compared to cascaded pipelines and also improves the overall translation quality [^1] Code is available at https://github.com/saikoneru/OmniFusion.
-OmniFusion: Simultaneous Multilingual Multimodal Translations via Modular Fusion
+Open-source text-only translation Large Language Models (LLMs) have rapidly improved in multilingual coverage and translation quality, but their unimodal design limits their applicability to multimodal translation.
+In speech translation (ST), they typically operate in cascaded pipelines with automatic speech recognition followed by translation, introducing additional latency particularly for simultaneous ST (SimulST) and they cannot exploit visual context for disambiguation in either speech or caption translation.
+In contrast, pretrained multimodal foundation models (MMFMs) provide strong cross-modal perception and reasoning but lack the multilingual depth and translation performance of specialized translation LLMs.
+We introduce OmniFusion, a unified architecture that combines a multimodal foundation model with a translation-specialized LLM through a lightweight gated fusion module.
+Built with Qwen2.5-Omni-7B as the MMFM and Seed-X-PPO-7B as the translation LLM, OmniFusion supports speech-to-text, speech-and-image-to-text, and text-and-image-to-text translation within a single architecture.
+Experiments show that OmniFusion effectively leverages audio and visual cues and reduces SimulST latency by one second compared to cascaded pipelines while maintaining competitive translation quality.
+[^1] Code is attached as a zip file.
+OmniFusion: Simultaneous Multilingual and Multimodal Translations via Modular Fusion
 Sai Koneru1, Matthias Huck2, and Jan Niehues1 1 Karlsruhe Institute of Technology 2 SAP SE, Dietmar-Hopp-Allee 16, 69190 Walldorf, Germany {sai.koneru, jan.niehues}@kit.edu {matthias.huck}@sap.com
 1 Introduction
-In recent years, research in natural language processing (NLP) has increasingly shifted from unimodal foundation models (Touvron et al., 2023; Grattafiori et al., 2024) to multimodal foundation models (MMFMs) (Abdin et al., 2024; Goel et al., 2025; Xu et al., 2025), as many tasks benefit from contextual information derived from multiple sources.
-Although unimodal systems can perform strongly in certain settings, further progress also requires models to be able to take advantage of additional contextual cues without degrading quality.
-One representative task that can benefit from such multimodal integration is speech translation (ST), where the goal is to translate spoken input in a source language into a target language.
-Over the years, advances in data scale, modeling techniques, and architectural design have led to substantial improvements in ST quality across many language pairs and domains (Ahmad et al., 2024; Agostinelli et al., 2025).
-However, the use of multimodal context—such as visual information from images or presentation slides accompanying speech—remains relatively underexplored (Liu and Niehues, 2024; Gaido et al., 2024a; Sinhamahapatra and Niehues, 2025).
+In recent years, NLP research has shifted from unimodal foundation models (Touvron et al., 2023; Grattafiori et al., 2024) to multimodal foundation models (MMFMs) (Abdin et al., 2024; Goel et al., 2025; Xu et al., 2025), as many tasks benefit from contextual information across sources.
+While unimodal systems can perform well in certain settings, further progress requires models that exploit additional cues without degrading quality.
+Translation is a representative task where multimodal integration is beneficial.
+In Speech Translation (ST), the goal is to convert spoken input in a source language into a target language (Ahmad et al., 2024; Agostinelli et al., 2025).
+Despite recent advances, leveraging visual context such as images or presentation slides accompanying speech remains underexplored (Liu and Niehues, 2024; Gaido et al., 2024a; Sinhamahapatra and Niehues, 2025).
+Similarly, caption translation involves short, image-linked text, where visual information can aid disambiguation and grounding.
 Figure 1: Architecture of OmniFusion with supported tasks and inference modes.
-Although MMFMs are capable of processing multiple modalities jointly, their multilingual translation capabilities are generally weaker than those of specialized translation large language models (LLMs) (Alves et al., 2024; Cheng et al., 2025).
-As a result, directly using MMFMs for ST often yields inferior translation quality compared to approaches that rely on strong text-based translation components (Appendix: Table 4).
-They also do not cover several languages compared to specialized translation LLMs.
-Consequently, a practical solution is to adopt a cascaded pipeline, where automatic speech recognition (ASR) is first performed, followed by machine translation (MT).
-This modular setup allows each component to be optimized independently and integrated during inference (Ahmad et al., 2024; Agostinelli et al., 2025; Koneru et al., 2025).
-However, since the translation model operates solely on text, it cannot exploit contextual information from the original speech or visual input.
-Moreover, the cascaded design introduces additional generation steps, leading to increased latency compared to end-to-end (E2E) models—an issue that is particularly critical for simultaneous ST (SimulST) applications (Agostinelli et al., 2025; Papi et al., 2025a).
-Therefore, we aim to develop an E2E MMFM that excels in multimodality, multilinguality and efficient for SimulST.
-This can be achieved by extending the capabilities of a translation LLM to handle additional modalities.
-Few prior works have explored adding audio or image inputs to multilingual LLMs (Ambilduke et al., 2025; Viveiros et al., 2025).
-However, these approaches typically rely on training from scratch or only using pretrained audio/vision encoders and do not leverage the higher-level representational capabilities of existing MMFMs.
-Furthermore, training is ineffective as the base LLM needs to learn the modality mapping from scratch.
-In this work, we propose leveraging pretrained MMFMs to introduce multimodal capabilities into a multilingual LLM with a E2E architecture.
-Rather than transferring only vision or audio encoders, we extract and utilize hidden-state representations from multiple layers of the MMFM to capture both the perceptual and reasoning components of multimodal understanding.
-Specifically, we adapt the translation LLM by incorporating hidden states from the first, middle, and last layers of the MMFM, thereby enabling the transfer of both low-level perception and high-level semantic understanding.
-To further strengthen cross-modal alignment, we introduce Optical Character Recognition (OCR) as a bridge task for the image modality, analogous to how ASR serves as an alignment mechanism for audio.
-This strategy allows the model to effectively leverage pretrained multimodal representations while receiving direct cross-modal supervision.
-We use Qwen Omni 2.5 7B (Xu et al., 2025) as the MMFM and SeedX 7B (Cheng et al., 2025) as the translation LLM to validate our approach on SimulST and multimodal MT tasks.
-We refer to the resulting fused model as OmniFusion.
-Our main contributions are summarized as follows:
-  • We propose a gated fusion approach that extracts multimodal embeddings from MMFMs, allowing translation LLMs to fully utilize the pretrained perceptual and interpretive capabilities of MMFMs.
-  • We show the benefit of using slides for context and achieve an improvement in latency of approximately $1$ second in the SimulST scenario compared to cascaded fine-tuned models while having better quality.
-(Figure 2).
-  • We observe a reduction in the major and critical errors for offline ST through self-cascading and effective use of the image context (Table 2).
-  • We achieve state-of-the-art performance in the CoMMuTE benchmark for image-text translation highlighting the ability of the model to exploit images (Table 3).
+Although MMFMs can process multimodal inputs, their multilingual translation performance is generally weaker than that of specialized translation LLMs (Alves et al., 2024; Cheng et al., 2025), leading to lower translation quality and limited language coverage.
+In ST, a common approach is a cascaded pipeline: automatic speech recognition (ASR) followed by machine translation (MT) (Ahmad et al., 2024; Agostinelli et al., 2025; Koneru et al., 2025).
+While this modular approach allows independent optimization, it prevents the translation model from leveraging multimodal context and adds latency, a limitation especially critical in simultaneous ST (SimulST) scenarios (Agostinelli et al., 2025; Papi et al., 2025a).
+These limitations motivate the development of an end-to-end multilingual and multimodal translation system supporting ST, image-aware ST, and caption translation.
+A unified model can exploit shared data across tasks and simplify deployment.
+Extending a translation LLM to handle additional modalities is one approach, but prior work (Ambilduke et al., 2025; Viveiros et al., 2025) often trains from scratch or relies solely on pretrained encoders, without fully leveraging the higher-level representations of existing MMFMs, making training inefficient.
+We propose leveraging pretrained MMFMs to introduce multimodal capabilities into a multilingual LLM within a unified E2E architecture.
+Our approach tightly couples the two models during training and employs a multi-task objective augmented with auxiliary tasks, including OCR and ASR with self-cascading, to enhance multimodal alignment and translation performance.
+Our contributions are threefold: (1) We present OmniFusion, an E2E model built using translation LLMs and pretrained MMFMs that jointly supports speech, text and image translation tasks.
+(2) We introduce a gated fusion architecture that selectively integrates representations from multiple MMFM layers into a translation LLM, enabling efficient training and inference for context-aware translations; and (3) We demonstrate that OmniFusion reduces SimulST latency by $\sim 1$ second (Figure 3), lowers major and critical error rates for offline ST (Table 1), and achieves state-of-the-art results on CoMMuTE (Table 2).
+OmniFusion demonstrates that MMFMs and translation LLMs can be effectively combined within a single architecture for multilingual multimodal translation.
 2 Approach
-We propose to fuse a MMFM with a task-specific translation LLM to leverage their complementary strengths and enable E2E training for improved latency.
-Furthermore, we aim to incorporate not only the perception capabilities of the MMFM but also its multimodal reasoning and understanding.
-First, we describe the overall architecture and training process of our approach, considering a general input and output sequence independent of the specific task.
-We formally define the information flow, training loss, and the fusion of MMFM layers.
-Next, we detail the training tasks used to align multimodal representations from both image and audio modalities to the translation LLM.
-Finally, we discuss the training prompts that enable a self-cascading inference mode, allowing the model to enhance translation quality at the cost of increased latency.
-An overview of the architecture and supported tasks is in Figure 1.
+We propose fusing a MMFM with a task-specific translation LLM to leverage their complementary strengths.
+The MMFM provides perceptual representations, while the translation LLM provides multilingual understanding and generation capabilities [^2] Using the MMFM as encoder does not provide significant computational overhead as decoders are primary latency bottleneck (Guta et al., 2025)..
+We first describe the overall architecture and training process, abstracting over task-specific inputs and outputs.
+Motivated by a probing experiment, we define the information flow, training loss, and fusion of MMFM layers.
+We then detail training tasks aligning image and audio modalities to the translation LLM and discuss prompts enabling self-cascading inference, which improves translation quality at higher latency.
+An overview is shown in Figure 1.
 2.1 Architecture
-Let us denote the specialized translation LLM as $\mathcal{M}_{\text{NMT}}$ and the MMFM as $\mathcal{M}_{\text{MMFM}}$, where both use a language model as their backbone.
-The $\mathcal{M}_{\text{NMT}}$ is trained for translation via supervised fine-tuning with the following loss function:
+Let $\mathcal{M}_{\text{NMT}}$ be the translation LLM and $\mathcal{M}_{\text{MMFM}}$ the MMFM, both with a language model backbone.
+$\mathcal{M}_{\text{NMT}}$ is trained via supervised fine-tuning with:
 $\mathcal{L}_{\text{Trans}}$ $=-\frac{1}{T_{y}}\sum_{t=1}^{T_{y}}\log P_{\theta}\big(y_{t}\mid y_{<t},\,X\big),$ (1)
@@
 where $X=(x_{1},\dots,x_{L_{x}})$ is the sequence of source tokens of length $L_{x}$, and $y=(y_{1},\dots,y_{T_{y}})$ is the sequence of target tokens of length $T_{y}$.
-For $\mathcal{M}_{\text{MMFM}}$, the image and audio modalities are first processed by modality-specific encoders and then projected to match the embedding dimension of $\mathcal{M}_{\text{MMFM}}$, denoted as $D_{\text{MMFM}}$.
-This projection allows the encoder outputs to be concatenated as a sequence of tokens, enabling the MMFM to operate similarly to a standard LLM, such as $\mathcal{M}_{\text{NMT}}$.
-To fuse both the perception and higher-level understanding of the MMFM, we hypothesize that lower layers predominantly capture perceptual features, while higher layers encode more abstract reasoning as computation progresses.
-Accordingly, we extract the first, middle, and last hidden layers from the MMFM, denoted as:
-$h^{(1)}_{t}$ $=\text{MMFM}_{\text{layer 1}}(x_{t}),$ (2)
-$h^{(\text{mid})}_{t}$ $=\text{MMFM}_{\text{layer mid}}(x_{t}),$
-$h^{(\text{last})}_{t}$ $=\text{MMFM}_{\text{layer last}}(x_{t}),$
-$h^{(1)}_{t},\,h^{(\text{mid})}_{t}$, $\,h^{(\text{last})}_{t}\in\mathbb{R}^{L_{m}\times D_{\text{MMFM}}},t=1,\dots,L_{m}$
+For $\mathcal{M}_{\text{MMFM}}$, image and audio inputs are encoded and projected to the MMFM embedding dimension $D_{\text{MMFM}}$, forming a token sequence compatible with the LLM.
+2.1.1 Probing MMFM Representations
+Figure 2: Average cosine similarity between hidden representations of adjacent layers, computed with Qwen-Omni 2.5 7B on the MCIF test set (en $\rightarrow$ de) under different modality conditions.
+To capture both perceptual and higher-level understanding, we hypothesize that the lower MMFM layers encode perceptual features and higher-level abstract reasoning.
+We validate this by computing cosine similarity between consecutive layer outputs, which is commonly used as a proxy for layer importance in model pruning literature.
+We compute with different modalities and visualize in Figure˜2.
+We see that the first and last layers exhibit the largest changes, while other layers provide small updates.
+Based on this, we propose using the first, middle, and last layers as inputs to generate multimodal representations as they capture the three regions [^3] We perform ablation using all layers in Table 3 and empirically validate this choice..
+2.2 Gated Fusion
+Accordingly, we extract these layers from the MMFM, denoted as:
+$h^{(k)}_{t}$ $=\text{MMFM}_{\text{layer }k}(x_{t}),\quad k\in\{1,\text{mid},\text{last}\},$ (2)
+$h^{(k)}_{t}$ $\in\mathbb{R}^{L_{m}\times D_{\text{MMFM}}},\quad t=1,\dots,L_{m}$
 where $L_{m}$ is the length of the MMFM hidden states.
-Directly concatenating these multi-dimensional representations into $\mathcal{M}_{\text{NMT}}$ poses three challenges.
-First, the sequence length triples, significantly increasing computational and memory requirements.
-Second, the relevance of each layer may vary depending on the task, necessitating selective weighting.
-Third, the embedding dimension of $\mathcal{M}_{\text{MMFM}}$ may not match that of $\mathcal{M}_{\text{NMT}}$.
-To overcome these challenges, we introduce a gated fusion layer that determines the contribution of each hidden state along the sequence length.
-Each hidden state is multiplied by its corresponding gate values, summed across layers, and then projected through an MLP to align the embedding dimension with that of the translation model:
+Directly concatenating these multi-dimensional MMFM representations into $\mathcal{M}_{\text{NMT}}$ is problematic because it triples the sequence length, layer relevance may vary by task, and embedding dimensions may not match.
+To address this, we use a gated fusion layer that weights each hidden state, sums across layers, and projects the result via an MLP to align with the translation model:
 $\begin{aligned} W_{\text{gate}}&\in\mathbb{R}^{3\times 3D_{\text{MMFM}}},\\ [g_{1},g_{\text{mid}},g_{\text{last}}]_{t}&=\text{softmax}\big(W_{\text{gate}}\cdot h^{(1)}_{t}\oplus h^{(\text{mid})}_{t}\oplus h^{(\text{last})}_{t}\big),\\ m^{\prime}_{t}&=g_{1}\odot h^{(1)}_{t}+g_{\text{mid}}\odot h^{(\text{mid})}_{t}+g_{\text{last}}\odot h^{(\text{last})}_{t},\\ m_{t}&=\text{MLP}(m^{\prime}_{t})\in\mathbb{R}^{L_{m}\times D_{\text{Trans}}},\quad t=1,\dots,L_{x}\end{aligned}$
-where $D_{\text{Trans}}$ is the embedding dimension of $\mathcal{M}_{\text{NMT}}$ and $m_{t}$ is the fused multimodal representation at time step $t$.
+where $\oplus$ is concatenation done per each token across layers, $D_{\text{Trans}}$ is the embedding dimension of $\mathcal{M}_{\text{NMT}}$ and $m_{t}$ is the fused multimodal representation at time step $t$.
 The fused multimodal embeddings $m_{t}$ can now be treated like additional token embeddings and concatenated with the $\mathcal{M}_{NMT}$ token embeddings:
@@
 2.3 Training
-We have shown how MMFM internal representations can be extracted and integrated into our downstream model.
-However, leveraging these representations effectively also requires a training process that explicitly aligns multimodal features so that the LLM can exploit them.
-The exact alignment procedure depends on the downstream tasks we expect the fused model to perform after fine-tuning.
-In this work, our goal is to obtain a translation model that can utilize both audio and image context.
-Therefore, we must make the translation LLM context-aware by training to exploit these representations.
-We consider $3$ translation tasks in this work and are defined as follows:
-  • Speech Translation (ST): Generate a translation in the target language given only audio in the source language.
-  • Speech-Image Translation (SIT): Generate a translation in the target language given source-language audio together with an image that provides additional context (e.g., a slide corresponding to the spoken segment).
-  • Text-Image Translation (TIT): Generate a translation in the target language given the source text and an aligned image, such as in caption translation scenarios.
-While we can train the model to perform translation directly, decomposing the task into intermediate steps can be beneficial.
-For example, translating English audio to German can be approached by first predicting ASR outputs and then producing the translation by continuing generation conditioned on the ASR hypothesis.
-This decomposition not only simplifies learning but also provides explicit alignment signals that help map audio representations into the LLM’s semantic space.
-When vision input is available—particularly for slide-based scenarios—we additionally introduce OCR as a bridge task before generating the final translation.
-In this case, the model must predict both OCR text and the final translation using the same audio–visual representation.
-This encourages alignment between the modalities and explicitly informs the model that the visual context is meaningful and should be utilized.
-During training, we stochastically decide—based on a specified probability for each data source—whether an example should be trained in self-cascading mode and along which modality the cascade should occur.
-Furthermore, we provide modality-specific prompt to the MMFM for handling different modality combination scenarios.
-Pseudocode for building the prompt for each task during training is shown in Algorithm˜1.
-We also freeze the MMFM to preserve its pretrained abilities while fine-tuning the translation LLM with LoRA (Hu et al., 2021).
-We use LoRA for computational benefits and not go Out-Of-Memory during fine-tuning.
+Effectively leveraging MMFM representations requires aligning multimodal features with the translation LLM’s token representations.
+We achieve this through multitask training that includes ST, ASR, and OCR objectives.
+2.3.1 Multi-task Learning
+We train the model on three primary multimodal translation tasks.
+In Speech Translation (ST), the model generates a target-language translation from source-language audio.
+In Speech–Image Translation (SIT), it translates audio together with an image providing additional context, such as a presentation slide aligned with the spoken segment.
+In Text–Image Translation (TIT), the model translates source text paired with an aligned image, as in caption translation scenarios.
+Training across these tasks exposes the model to different modality combinations, enabling effective alignment and exploitation of multimodal representations.
+2.3.2 Self-Cascading (ASR and OCR)
+In addition to multi-task learning, we adopt self-cascading, which decomposes translation into intermediate steps.
+For English-to-German ST, the model first generates the ASR transcription, then produces the final translation conditioned on it, providing explicit alignment between audio representations and the LLM’s semantic space.
+For visual inputs, OCR text serves as an intermediate step before translation, aligning visual and textual representations.
+During training, self-cascading is applied stochastically based on task and modality, with modality-specific prompts (Algorithm˜1).
+We freeze the MMFM and fine-tune the translation LLM with LoRA (Hu et al., 2021), reducing compute costs while retaining pretrained multimodal capabilities.
+We also find that adding LoRA to the MMFM degrades performance due to catastrophic forgetting.
 3 Experimental Setup
 3.1 Training Data
-We aim to train the fused model to perform translation tasks across both speech and image modalities. An overview of the data sources is provided in Table˜1.
-For ST, we use subsets of Europarl-ST (Iranzo-Sánchez et al., 2020) and Covost-2 (Wang et al., 2020), sampling across multiple target languages.
-We limit training data to these subsets due to computational constraints and to demonstrate that full datasets are not strictly necessary.
-To enable the model to exploit images when translating audio, we use the M3AV corpus (Chen et al., 2024), which contains English scientific talks paired with corresponding slides.
-The slides also include Paddle-OCR (Cui et al., 2025) predictions, allowing the model to be trained either to predict translations directly or via an intermediate OCR step for better alignment.
-Since M3AV has limited language coverage, we augment the data by synthetically translating ASR transcripts into additional target languages (M3AV-Aug).
-This enables the model to handle ST with slides and supports the self-cascade mode, performing ASR first and then translation.
-Finally, for TIT, we employ the Multi30k dataset (Elliott et al., 2016), which contains images paired with captions in multiple languages.
-This task encourages the model to go beyond OCR and to leverage visual content, including objects and scene understanding, for translation.
+We train the fused model to perform translation across speech and image modalities.
+An overview of data sources is provided in Appendix Table˜6.
+For ST, we use subsets of Europarl-ST (Iranzo-Sánchez et al., 2020) and Covost-2 (Wang et al., 2020), sampling multiple target languages.
+Training is limited to these subsets to reduce computational costs.
+To exploit images in audio translation, we use the M3AV corpus (Chen et al., 2024), containing English scientific talks paired with slides.
+The slides include Paddle-OCR (Cui et al., 2025) predictions, allowing training either for direct translation or via an intermediate OCR step for better alignment.
+Since M3AV has limited language coverage, we augment ASR transcripts with synthetic translations (M3AV-Aug), enabling ST with slides and supporting self-cascade training.
+For TIT, we use Multi30k (Elliott et al., 2016), which pairs images with multilingual captions.
+This task encourages leveraging visual content beyond OCR, including objects and scene understanding, to improve translation.
 3.2 Models
-Several MMFMs have been developed rapidly in recent years, showing significant improvements in both quality and efficiency (Abdin et al., 2024; Team et al., 2025).
-In our scenario, we require a model capable of processing both audio and image modalities, while also supporting cross-modal abilities.
-Accordingly, we select Qwen Omni 2.5B [^2] Qwen/Qwen2 Audio.5-Omni-7B (Xu et al., 2025), due to its strong multimodal capabilities and the potential to incorporate video in future experiments.
-For the translation LLM, we choose the recently proposed SeedX PPO 7B [^3] ByteDance-Seed/Seed-X-PPO-7B (Cheng et al., 2025), which provides broad language coverage and strong translation performance compared to closed-source alternatives.
-We use the training data mentioned above and denote the fine-tuned model as OmniFusion for the rest of the paper.
-While we report results using these specific models, it is important to note that our approach is not limited to these base LLMs.
-The proposed fusion strategy does not require massive pre-training or fine-tuning resources and can be adapted in more modest computing environments, such as university clusters.
-The training hyper-parameters are specified in Table˜10.
-For validation, we used Covost-2 validation data and select the checkpoint with the least loss in 20k steps.
-Dataset # Examples Language Modality Task
-M3AV 122k en Speech, Image ASR, OCR
-M3AV - Aug 122k
-zh, fr, de, it,
-ja, ko, pt, ru,
-es, vi, ar, cs,
-hr, da, nl, fi,
-hu, id, ms, nb,
-no, pl, ro, tr
-Speech, Image ST, OCR
-Europarl 260k*
-de, es, fr, it, nl,
-pl, pt, ro
-Speech ST
-Covost2 de, ja, tr, zh Speech ST
-Coco-Captions 87k cs, de, fr Image TIT
-Table 1: Overview of data sources collected for E2E training.
-* We sample together 260k from the concatenated data uniformly across the language pairs.
-All the speech and text data are only available with English as the source language.
-Language code mapping is presented in Table˜10.
+Several MMFMs have recently shown strong gains in quality and efficiency (Abdin et al., 2024; Team et al., 2025).
+As our setting requires processing both audio and image modalities with cross-modal reasoning, we use Qwen Omni 2.5 7B [^4] Qwen/Qwen2.5-Omni-7B (Xu et al., 2025).
+For translation, we adopt SeedX PPO 7B [^5] ByteDance-Seed/Seed-X-PPO-7B (Cheng et al., 2025), which provides broad language coverage and strong translation performance compared to closed-source alternatives.
+Using the training data described above, we refer to the resulting fine-tuned model as OmniFusion.
+While we report results with these models, the approach is general and applicable to other base LLMs.
+Training hyperparameters are listed in Table˜14.
+Validation uses Covost-2 dev data, selecting the checkpoint with the lowest loss after 20k steps.
 3.3 Evaluation: Data & Metrics
-For evaluating the performance of multimodal Simul-ST systems, we require aligned audio-visual translation data.
-Additionally, we avoid using M3AV for testing, as it overlaps with our training data, monolingual and we aim to measure the model’s robustness on unseen domains.
-Consequently, we use the MCIF test set (Papi et al., 2025b), which contains video and audio recordings of ACL talks in three target languages from English.
-We report latency with Average Lagging (Ma et al., 2020) and translation quality results on this dataset.
-For TIT, we focus on test cases where the image is crucial for disambiguating the source sentence.
-For this, we use the CoMMuTE dataset (Futeral et al., 2023, 2025), which provides a small but multilingual targeted TIT test set.
-We measure translation quality using XCOMET-XL (Guerreiro et al., 2024) for ST scenarios, as it offers high-quality evaluation and allows analysis of minor, major, and critical errors.
-For TIT, we use the COMET [^4] Unbabel/wmt22-comet-da model (Rei et al., 2022) directly, without providing error spans, to evaluate aligned references.
+Evaluating multimodal SimulST requires aligned audio–visual translation data.
+We exclude M3AV from testing due to training overlap and its monolingual nature, and instead assess robustness on unseen domains using the MCIF test set (Papi et al., 2025b), which contains ACL talk recordings translated from English into three target languages, and FLEURS (Conneau et al., 2023) for multilingual coverage.
+Latency is reported using Average Lagging (Ma et al., 2020), and translation quality is measured on these datasets.
+For TIT, we target cases where visual context is crucial for disambiguation, using the CoMMuTE benchmark (Futeral et al., 2023, 2025).
+Translation quality is evaluated with XCOMET-XL (Guerreiro et al., 2024) for ST, enabling detailed error analysis, and with COMET [^6] Unbabel/wmt22-comet-da (Rei et al., 2022) for TIT.
 4 Results
-Our primary motivation for the proposed approach is to build an E2E ST system that is context-aware, efficient for SimulST and enabled to exploit additional image context.
-To validate this, we first evaluate the system in simultaneous scenarios by comparing latency and quality.
-Next, we report quality in offline scenarios and conduct an error analysis to contrast the E2E system with traditional cascaded systems fine-tuned on the same data.
-We then assess performance in TIT scenarios, benchmarking against other off-the-shelf multimodal translation models.
-Finally, we perform an ablation study to analyze the contributions of individual layers in the MMFM and present the key findings.
-We also validate the proposed approach for a different model combination (Qwen-Audio (Chu et al., 2023) and Tower-Instruct (Alves et al., 2024)), analyze training with different layers and post-editing in our preliminary experiments and report the results in the Appendixes Table˜4 and table˜5.
+Our goal is to develop a high-quality multimodal translation system with low-latency end-to-end SimulST that effectively leverages image context.
+We evaluate latency–quality trade-offs in simultaneous settings, compare offline translation quality and errors against cascaded models trained on the same data, assess text-and-image translation performance against multimodal baselines, and present ablations on MMFM layers and fusion mechanisms.
+We also report ablations on training strategies to disentangle architectural and training effects (Appendix Table˜9).
+Additionally, we include preliminary experiments with alternative model pairings and layer selections (Qwen-Audio (Chu et al., 2023) and Tower-Instruct (Alves et al., 2024)) as well as post-editing results (Appendix Tables˜5 and 8).
+We observe similar findings and do not explore additional model combinations due to computational constraints.
 (a) en $\rightarrow$ de
@@
 4.1 Simultaneous Speech Translation
-A SimulST system must generate translations incrementally as the source audio arrives, rather than waiting for the entire utterance to finish.
-The decision of whether to commit a token or read more audio is referred to as the policy.
-Since our E2E system does not natively support simultaneous translation, we adopt the widely used Local Agreement policy (Liu et al., 2020), which commits the longest common prefix observed across consecutive predictions.
-Concretely, at each step we process the audio in fixed chunk sizes (a hyperparameter) and irreversibly emit the overlapping prefix between the current and previous hypotheses.
-After committing, subsequent tokens are force-decoded conditioned on the partial translation committed so far.
-To fairly evaluate the proposed approach, we compare our E2E model OmniFusion against a strong cascaded counterpart built from the same pretrained components.
-We therefore fine-tune this cascaded baseline on the same data using LoRA adapters (Omni + SeedX LoRA FT).
-The effect of fine-tuning is discussed in more detail in Section˜4.2.
-We experiment with chunk sizes of $[1,1.5,2,2.5,3]$ seconds, run inference both with and without images, and report computation-aware and computation-unaware AL together with XCOMET-XL scores for the $en\rightarrow de$ and $en\rightarrow it$ language pairs in Figure˜2.
-The results reveal several key insights.
-First, as expected, the offline ST systems (with images) achieve the highest quality in both language pairs due to access to full context.
-Next, OmniFusion is consistently around 1 second faster than its cascaded counterpart across all chunk sizes while preserving comparable quality in the audio-only setting.
-When images are provided, OmniFusion delivers both the lowest latency and the highest quality, demonstrating that visual context is substantially more beneficial in the E2E model than in the cascaded baseline allowing it to commit high quality stable outputs earlier.
-Finally, we also see that in computation-aware settings, adding image causes delay due to processing and causes overhead than only-audio creating a trade-off between quality and latency.
-We attribute the advantage of OmniFusion to two factors: (1) our fine-tuning strategy, which includes an OCR objective that tightly aligns image and audio representations, and (2) the inherent efficiency of the E2E architecture, which requires fewer decoding steps overall.
-In contrast, the cascaded pipeline with Local Agreement must wait for stable ASR hypotheses before translation can even begin, making it slower than the E2E approach.
-In conclusion, OmniFusion achieves better latency, leverages image context effectively, and has higher quality in SimulST scenarios, making it particularly well-suited for real-world applications such as live presentations and meetings where visual information is readily available.
+A SimulST system must translate incrementally as audio arrives, requiring a policy to decide when to read or commit output.
+Since our E2E model does not natively support simultaneous decoding, we adopt the Local Agreement policy (Liu et al., 2020), which commits the longest common prefix between consecutive hypotheses.
+Audio is processed in fixed-size chunks, and committed tokens are force-decoded in subsequent steps.
+For a fair comparison, we evaluate OmniFusion against a strong cascaded baseline built from the same pretrained components and fine-tuned on the same data using LoRA adapters (Omni + SeedX).
+We experiment with chunk sizes of $[1,1.5,2,2.5,3]$ seconds, with and without images.
+As computational time in SimulST is critical, we report both computation-aware and -unaware Average Lagging along with XCOMET-XL scores for $en\rightarrow de$ and $en\rightarrow it$ in Figure˜3.
+Offline ST with images yields the highest quality, as expected.
+In simultaneous settings, OmniFusion is consistently about 1 second faster than the cascaded system across all chunk sizes, with comparable or slightly lower quality in the audio-only condition.
+When images are provided, OmniFusion achieves both the lowest latency and highest quality, indicating that visual context is more effectively exploited in the E2E model, enabling earlier and more stable commitments.
+Image processing introduces additional delay, highlighting a latency–quality trade-off.
+We hypothesize that OmniFusion’s advantage stems from (i) our fine-tuning strategy, including an OCR objective that aligns audio and visual representations (Table˜9 shows performance degradations without multitask objectives) and (ii) the efficiency of the end-to-end architecture, which prevents the delayed onset of translation inherent in cascaded pipelines.
+Combined with Local Agreement decoding, OmniFusion achieves improved latency–quality trade-offs compared to cascaded systems.
 4.2 Offline Speech Translation
 Category Config Image XCOMET-XL Minor Major Critical Total
-Cascade Systems (ASR → MT)
-Omni + SeedX
-No FT 85.88 1212.3 773.0 74.3 2059.7
+Omni FT (Direct ST) LoRA FT 82.30 1079.0 852.0 90.3 2021.3
+Cascade Systems (ASR → MT) Omni + SeedX No FT 85.88 1212.3 773.0 74.3 2059.7
 No FT 85.94 1191.3 759.3 74.7 2025.3
@@
 LoRA FT 86.59 1202.3 766.3 55.7 2024.3
-End-to-End (Direct ST)
-OmniFusion
-Mid Fusion 83.57 1114.0 1188.3 72.3 2374.7
+End-to-End (Direct ST) OmniFusion Mid Fusion 83.57 1114.0 1188.3 72.3 2374.7
 Mid Fusion 85.98 1256.3 749.0 61.7 2067.0
@@
 Gated Fusion 86.24 1255.7 736.0 57.0 2048.7
-Gated Fusion + Self-Casade 86.57 1231.0 719.0 55.3 2005.3
+Gated Fusion + Self-Cascade 86.57 1231.0 719.0 55.3 2005.3
 […]
-Although SimulST is the primary focus of this work, we also evaluate the systems in the offline scenario, where latency is not a concern and the goal is to maximize translation quality.
-In this setting, we can leverage the self-cascade inference mode described in Section˜2.2 to further boost performance.
-To measure the contributions of fine-tuning and visual context, we first evaluate the cascaded baseline in four configurations: base vs.
-fine-tuned (LoRA FT) models, each tested with and without images.
-For the E2E OmniFusion model, we additionally train a variant that uses only the middle-layer fusion without gating to assess the need for the gating module.
-We then compare all E2E configurations (mid-only fusion and full gated fusion) both with and without images, and with/without self-cascade inference.
-Results are averaged across three target languages $en\rightarrow\{de,it,zh\}$ and reported in Table˜2 using XCOMET-XL scores.
-We further include the detailed XCOMET-XL error categories, hypothesizing that E2E models may sacrifice minor fluency for substantially fewer major and critical errors.
-The results yield several insights.
-In the cascaded pipeline, LoRA fine-tuning yields consistent gains (e.g., 85.94 $\rightarrow$ 86.59 with images), whereas adding images provides only marginal benefit.
-In contrast, the E2E OmniFusion model benefits substantially from both gated fusion (outperforming mid-only fusion) and image context showing the need for gating module and additional contextual information.
-Although the direct E2E translation score remains slightly below the best cascaded system, enabling self-cascade inference closes this gap almost entirely (cascaded LoRA FT: 86.59 vs.
-E2E gated fusion + self-cascade: 86.57).
-More importantly, error analysis reveals that our best E2E configuration produces the fewest critical and major errors (e.g., 719 vs.
-751.3 for the strongest cascaded baseline).
-These findings demonstrate that OmniFusion, when equipped with image context and self-cascade inference, matches fine-tuned cascaded system in offline ST quality while offering significantly fewer severe translation errors
-4.3 Text-Image Translation
 Model Config en $\rightarrow$ ar en $\rightarrow$ cs en $\rightarrow$ de en $\rightarrow$ fr en $\rightarrow$ ru en $\rightarrow$ zh
@@
 SeedX 7B - 80.39 83.56 82.38 80.31 82.13 83.25
-ZeroMMT 3.3B
-(Futeral et al., 2025)
-- 80.64 86.69 83.54 83.40 83.87 77.85
-TowerVision 9B
-(Viveiros et al., 2025)
-- 70.85 86.62 84.97 84.29 85.81 85.24
+ZeroMMT 3.3B - 80.64 86.69 83.54 83.40 83.87 77.85
+TowerVision 9B - 70.85 86.62 84.97 84.29 85.81 85.24
 OmniFusion 14B Mid Fusion 80.98 87.49 85.14 84.97 85.69 85.15
 OmniFusion 14B Gated Fusion 80.81 87.58 84.88 84.28 84.62 84.34
-Table 3: COMET scores ($\uparrow$) on the CoMMuTE dataset.
-Statistically significant top-performing results per language pair are highlighted in bold.
-‘Config‘ distinguishes the two OmniFusion variants (mid-layer fusion vs.
-gated fusion).
-Apart from ST, translating captions also relies on image context, where cascaded systems fail due to their inability for disambiguation.
-To demonstrate this, we evaluate several models on the CoMMuTE test set across multiple language pairs and report COMET scores in Table˜3.
-First, the audio-only Omni 2.5B exhibits the weakest multilingual performance, which reinforces our motivation for integrating it with a strong translation backbone.
-SeedX 7B obtains reasonable scores yet consistently trails all vision-capable models, confirming its limitation in the absence of image context.
-Among vision-aware systems, both Mid Fusion and Gated Fusion variants of our OmniFusion 14B achieve state-of-the-art or near-SOTA COMET scores, matching or surpassing vision-only models such as ZeroMMT 3.3B (Futeral et al., 2025) and TowerVision 9B (Viveiros et al., 2025) on most language directions.
-Interestingly, and in contrast to ST, the Mid Fusion variant slightly outperforms Gated Fusion on CoMMuTE.
-This suggests that, when the image serves primarily as supplementary context (as in caption translation) rather than the main information source (as in audio for spoken ST), using visual features at the middle layer alone may be sufficient, and the additional flexibility of gating offers no further advantage.
-Importantly, unlike TowerVision which requires massive multimodal pretraining from scratch, OmniFusion can be built on top of existing strong pretrained components.
-This makes our approach far more data- and training-efficient without compromising quality, by effectively exploiting the rich image–text alignment knowledge already present in the MMFM pretrained weights.
-These results underline the strength of our E2E design on visually grounded translation tasks where image context is essential.
-4.4 Contribution of Layers
+Table 2: COMET scores ($\uparrow$) on the CoMMuTE dataset. Statistically significant top-performing results per language pair are highlighted in bold. Qualitative example can be found in Table˜16
+Although SimulST is our primary focus, we also evaluate all systems in an offline setting where latency is unconstrained and translation quality is maximized.
+In this scenario, we additionally leverage the self-cascade inference mode (Section˜2.3) to further improve performance.
+To isolate the effects of fine-tuning and visual context, we evaluate the cascaded baseline in five configurations: base MMFM fine-tuned for ST [^7] Even if it is competitive, it is insufficient due to limited language coverage compared to SeedX, and base vs.
+LoRA cascade fine-tuned with and without images.
+For OmniFusion, we train an additional mid-only fusion variant (without gating) (We find using the middle layer performs best in our preliminary experiments from Table 5) to assess the contribution of the gating module.
+We compare all E2E configurations with and without images and self-cascade inference.
+Results are averaged over $en\rightarrow{de,it,zh}$ and reported using XCOMET-XL in Table˜1, including fine-grained error categories.
+Additionally, inference times are reported in Table 7, although latency is not a primary constraint.
+First, directly fine-tuning the MMFM for speech translation performs substantially worse than OmniFusion (82.30 vs 83.98 XCOMET), confirming combining is more effective than adapting the MMFM alone.
+In the cascaded pipeline, LoRA fine-tuning provides consistent gains (e.g., 85.94 $\rightarrow$ 86.59 with images), while images offer only marginal improvements.
+In contrast, the E2E OmniFusion model benefits substantially from both gated fusion (outperforming mid-only fusion) and image context, highlighting the importance of gating and additional contextual information.
+Although direct E2E translation slightly trails the best cascaded system, enabling self-cascade inference nearly closes the gap (cascaded LoRA FT: 86.59 vs.
+E2E gated fusion + self-cascade: 86.57).
+More importantly, error analysis shows that our best E2E configuration produces fewer critical and major errors (e.g., 719 vs.
+751.3 major errors for the best cascaded system).
+These results demonstrate that OmniFusion with image context and self-cascade inference matches fine-tuned cascaded systems in offline ST quality while producing fewer severe translation errors.
+4.3 Text-Image Translation
+Caption translation relies heavily on visual context, where cascaded systems often fail due to limited disambiguation.
+To evaluate OmniFusion’s ability to leverage images, we report COMET scores on the CoMMuTE test set across multiple language pairs in Table˜2.
+Omni 2.5B exhibits the weakest multilingual performance, motivating its fusion with a stronger translation backbone.
+SeedX 7B achieves reasonable scores but consistently underperforms vision-aware models, highlighting its limitations without image context.
+Among multimodal systems, both Mid Fusion and Gated Fusion variants of OmniFusion 14B reach near state-of-the-art performance, matching or surpassing vision-only models such as ZeroMMT 3.3B (Futeral et al., 2025) and TowerVision 9B (Viveiros et al., 2025) on most language directions.
+Unlike in ST, Mid Fusion slightly outperforms Gated Fusion on CoMMuTE, suggesting that when images provide auxiliary rather than primary information, mid-layer visual integration is sufficient.
+Notably, unlike TowerVision, which relies on large-scale multimodal pretraining from scratch, OmniFusion builds on strong pretrained components, offering a more data- and training-efficient approach without sacrificing quality.
+These results highlight the context-awareness of our E2E design for visually grounded translation.
+4.4 Layer Contribution - Training
+Model Mode Fusion Layers en $\rightarrow$ de cn $\rightarrow$ de jp $\rightarrow$ ru fr $\rightarrow$ ko ur $\rightarrow$ cn
+Chrf Comet-22 Chrf Comet-22 Chrf Comet-22 Chrf Comet-22 Chrf Comet-22
+Tower-Spire Direct 56.9 82.54 N/A N/A N/A N/A N/A N/A N/A N/A
+Tower-Spire Self-Cascade 61.9 85.02 N/A N/A N/A N/A N/A N/A N/A N/A
+SeamlessM4T v2 Direct 62.4 85.60 35.8 70.22 36.9 71.77 20.8 74.12 13.5 61.11
+SeamlessM4T v2 Self-Cascade 63.3 85.81 36.9 71.34 38.2 69.85 21.0 75.22 13.5 63.75
+Our Models - OmniFusion
+Attention Direct All 60.9 84.02 37.4 72.96 26.6 65.62 21.2 75.36 4.3 43.66
+Attention Direct 3 61.8 86.95 39.4 74.86 27.6 68.81 21.9 78.94 5.5 46.78
+Gated Direct All 62.3 85.37 39.6 75.91 28.8 69.33 21.8 77.50 5.0 45.78
+Gated Direct 3 64.0 87.12 44.5 81.82 37.2 81.95 25.5 82.74 12.5 65.60
+Table 3: Ablation study of fusion architecture with non-english zero-shot source languages. Tower-Spire only supports English audio.
+We use only the first, middle, and last layers in the gated fusion.
+To validate this choice, we perform ablation experiments where all layers of the MMFM are used within the gating module.
+In addition, we explore an attention-based fusion approach, where a learned query token attends over all layer hidden states to produce new contextualized representations.
+Moreover, we compare our method with other state-of-the-art models and report the results in Table 3 on the FLEURS test set to account for multilinguality.
+First, we observe that the gated variants achieve the best performance across all language pairs, including challenging ones such as Urdu to Chinese.
+Next, using all layers in the gating module leads to performance degradation, which becomes even more pronounced with attention-based fusion.
+We attribute this primarily to the increased noise introduced when incorporating all layers.
+Further, attention-based fusion strategies fail compared with gating due to training/test mismatch, especially for non english source languages.
+In comparison, our proposed approach introduces minimal number of parameters, avoiding reliance on large datasets while still achieving competitive performance.
+4.5 Layer Contribution - Inference
 Figure 4: Analysis of the layer contribution from the MMFM in the gating module across modalities.
-We presented a gating module in Section˜2.1 for fusing information across layers of the MMFM.
-To better understand its behavior, we conduct an ablation study by examining the gate values assigned to each MMFM layer during fusion.
-Specifically, we compute the average gate value across the three layers (first, middle, last), noting that the gate weights for each timestep sum to 1.
-We run inference on 100 examples from the MCIF dataset—both speech-only and speech+image conditions—and on the CoMMuTE dataset for the image-only modality.
-The resulting average gate values are shown in Figure˜3.
-We observe that the first and middle layers receive consistently higher contributions, while the last layer contributes only marginally.
-This suggests that the majority of transferable multimodal processing occurs in the earlier half of the MMFM.
-Although the weak contribution of the last layer is not entirely clear, we hypothesize that these representations may be more language-specific, whereas the middle layers capture more language-agnostic features (Liu and Niehues, 2025).
+We introduced a gating module (Section˜2.1) to fuse information from multiple MMFM layers.
+To analyze its behavior, we perform an ablation study by examining the average gate values assigned to the first, middle, and last MMFM layers, noting that gate weights sum to 1 at each timestep.
+We run inference on 100 MCIF examples under speech-only and speech+image conditions, and on CoMMuTE for image-only input.
+The resulting average gate values are shown in Figure˜4.
+Across settings, first and middle layers receive substantially higher weights, while the last layer contributes marginally.
+This indicates that most transferable multimodal information is captured in earlier MMFM layers.
+Inference-time Gate Force Speech Speech + Image
+First 55.22 63.02
+Middle Halluc. + ASR Halluc. + ASR
+Last EOS EOS
+89.90 90.05
+Table 4: Inference time control of the gating with forcing first, middle or last layer. We report XCOMET-XL scores on the MCIF en $\rightarrow$ de test set.
+Next, we conduct another ablation by performing inference-time control in the gating module.
+We ignore the gate’s predictions and instead force the model to use only the first, middle, or last layer, and report the results in Table 4.
+Forcing the first layer degrades translation quality but does not collapse the system and even improves performance when images are present.
+In contrast, forcing the middle or last layer leads to failure modes.
+With the middle layer, the system often generates the transcript along with additional noise, while forcing the last layer causes the model to predict only the eos token.
+We hypothesize that the first layer provides reasonably aligned multimodal embeddings, whereas the middle layer contains more processed representations that become ineffective due to a training–test mismatch.
+Notably, training with only a single layer still produces workable systems, although using the last layer yields the lowest performance and weakest zero-shot transfer, as shown in Table 5.
 5 Related Work
-Multimodal MT: Translating captions has been widely explored in the MT field over the last decade (Specia et al., 2016; Elliott et al., 2017; Barrault et al., 2018).
-Several works focus on enabling NMT models to process images, either by training models from scratch (Vijayan et al., 2024) or by augmenting large pretrained models such as NLLB (Costa-Jussà et al., 2022; Futeral et al., 2025).
-Most recently, and in parallel to our work, Viveiros et al.
+Multimodal MT: Translating captions has been widely explored over the past decade (Specia et al., 2016; Elliott et al., 2017; Barrault et al., 2018).
+Several works extend NMT models to process images, either by training from scratch (Vijayan et al., 2024) or by augmenting pretrained models such as NLLB (Costa-Jussà et al., 2022; Futeral et al., 2025).
+Recently, and in parallel with our work, Viveiros et al.
 (2025) extend Tower with image and video modalities through a computationally expensive multi-stage training pipeline.
-While these efforts focus primarily on vision, an other line of direction is adding audio to translation LLMs (Ambilduke et al., 2025).
-A smaller line of work tackles the combination of vision and audio, such as Sinhamahapatra and Niehues (2025), which exploits slides and speech for ASR.
-Building on this trend, there is also increasing interest in translating audio together with video (Anwar et al., 2023; Choi et al., 2024).
-In contrast to these works, our approach focuses on jointly using images and audio specifically for ST, while remaining efficient in both training and inference.
-Simultaneous ST: Many real-world scenarios require real-time ST without waiting for the full source utterance.
-However, most research focuses on training models in offline settings and only later adapting them to SimulST.
+Beyond vision, another line of research explores incorporating audio into translation LLMs (Ambilduke et al., 2025).
+A smaller body of work investigates the joint use of vision and audio, such as Sinhamahapatra and Niehues (2025), which leverages slides and speech for ASR.
+Recent studies also explore translating with video (Anwar et al., 2023; Choi et al., 2024).
+In contrast, our approach focuses on jointly leveraging images and audio while remaining efficient for data and latency.
+Simultaneous ST: Real-world scenarios require real-time ST without waiting for the full utterance.
+However, most research trains models in offline settings and later adapts them to SimulST.
 Such adaptation typically relies on predefined or learned translation policies, such as wait-k (Ma et al., 2019; Elbayad et al., 2020), which assumes a fixed token delay before committing to output.
-Other approaches include Local Agreement (Liu et al., 2020), which selects the most stable prefix across consecutive read steps or attention-based heuristics (Papi et al., 2023).
+Other approaches include Local Agreement (Liu et al., 2020), which selects the most stable prefix across consecutive read steps, and attention-based heuristics (Papi et al., 2023).
 More recently, several works explore leveraging LLMs for SimulST policy learning (Koshkin et al., 2024b, a; Guo et al., 2025; Fu et al., 2025).
-In our work, we use Local Agreement to validate the effectiveness of our approach and to compare against cascaded baselines.
-However, our method is compatible in theory with other SimulST policies, and integrating more advanced or LLM-driven strategies is a promising direction for future work.
-Modality Fusion: With the rapid development of multimodal foundation models, a key challenge is how to effectively integrate speech and image modalities into text-centric LLMs.
-Some approaches (Abdin et al., 2024; Xu et al., 2025; Team et al., 2025) process each modality with separate encoders and merge them later within the LLM backbone.
+In this work, we adopt Local Agreement to evaluate our approach and compare against cascaded baselines.
+Our method is, however, compatible with other SimulST policies given its E2E nature.
+Modality Fusion: A key challenge is how to effectively integrate speech and image modalities into text-centric LLMs.
+Some approaches (Abdin et al., 2024; Xu et al., 2025; Team et al., 2025) process each modality with separate encoders and merge them within the LLM.
 In contrast, models such as (Ye et al., 2025) enable early interaction between vision and audio modalities, facilitating better cross-modal alignment.
-Beyond the timing of modality interaction, ongoing research also investigates the design of modality projectors (Verdini et al., 2024; Gaido et al., 2024b) and the use of continuous versus discrete tokens for non-text modalities (Zhan et al., 2024; Li et al., 2025), balancing the trade-offs between latency and translation quality.
+Beyond the timing of such interaction, recent works also investigates the design of modality projectors (Verdini et al., 2024; Gaido et al., 2024b) and the use of discrete tokens (Zhan et al., 2024; Li et al., 2025), balancing between latency and translation quality.
 6 Conclusion
-We proposed a novel approach to fuse a MMFM with a translation LLM, and validated its effectiveness across three key scenarios.
-First, we showed that the fused model achieves faster inference while maintaining strong translation quality in SimulST settings, even when only partial source audio is available.
-Second, the model produces fewer major and critical translation errors, demonstrating more reliable linguistic accuracy.
-Third, we found that the model effectively leverages visual context for both speech and text translation, leading to consistent gains in multimodal settings.
-Finally, our layer-level analysis reveals that the first and middle MMFM layers contribute most to capturing semantic multimodal information, providing insight into where meaningful cross-modal representations emerge.
-For future work, our fusion framework can naturally be extended to additional modalities.
-Since Omni already supports video processing, incorporating video embeddings would enable richer context for translation.
-Moreover, the fusion strategy is not limited to models for multimodal inputs but future extensions could target multimodal generation, allowing the system to produce not only text but also speech, images, or video as output.
+We propose a novel approach for fusing an MMFM with a translation LLM to enable high-quality, multilingual, multimodal, and simultaneous translation.
+We validate its effectiveness across three scenarios.
+First, the fused model achieves lower latency while maintaining strong translation quality in SimulST, even with partial source audio.
+Second, it produces fewer major and critical errors.
+Third, it effectively exploits visual context for both speech and text translation, yielding consistent gains in multimodal settings.
+Our layer-level analysis further shows that the first and middle MMFM layers contribute most to semantic multimodal representations.
 Limitations
-A primary limitation of our work is that OmniFusion is trained with fixed task-specific prompts across the three supported translation settings, and therefore does not exhibit general instruction-following capabilities.
+Our training data is primarily English-centric, which may limit performance for other source languages.
+Additionally, while we evaluate multimodal translation on established benchmarks, broader human evaluation would help better assess the model’s ability to utilize visual context.
+Future work will explore multilingual multimodal training and larger-scale human evaluation Another limitation of our work is that OmniFusion is trained with fixed task-specific prompts across the three supported translation settings, and therefore does not exhibit general instruction-following capabilities.
 It remains unclear whether our fusion strategy can be extended to support more flexible, instruction-driven multimodal behavior.
-In addition, most of our experiments use English as the source language.
-Although we report zero-shot results in Table 4, we do not evaluate the OmniFusion model with vision inputs in multilingual zero-shot conditions.
 Appendix A Appendix
@@
 Underlined langauges are not seen during the fusing process as source/target.
+Dataset # Examples Language Modality Task
+M3AV 122k en Speech, Image ASR, OCR
+M3AV - Aug 122k
+zh, fr, de, it,
+ja, ko, pt, ru,
+es, vi, ar, cs,
+hr, da, nl, fi,
+hu, id, ms, nb,
+no, pl, ro, tr
+Speech, Image ST, OCR
+Europarl 260k*
+de, es, fr, it, nl,
+pl, pt, ro
+Speech ST
+Covost2 de, ja, tr, zh Speech ST
+Coco-Captions 87k cs, de, fr Image TIT
+Table 6: Overview of data sources collected for E2E training.
+* We sample together 260k from the concatenated data uniformly across the language pairs.
+All the speech and text data are only available with English as the source language.
+Language code mapping is presented in Table˜14.
+Model Time(s) Image XCOMET-XL
+Omni + SeedX FT 2.96 w/o 89.75
+OmniFusion 1.98 w/o 88.09
+OmniFusion SC 3.78 w/o 88.44
+Omni + SeedX FT 3.85 w/ 90.18
+OmniFusion 3.15 w/ 89.90
+OmniFusion SC 4.82 w/ 90.05
+Table 7: Quality vs latency tradeoffs in the offline setting for English $\rightarrow$ German on the MCIF test set. SC indicates Self-Cascade mode.
 Model Mode Hypothesis Comet-22 ($\uparrow$)
@@
 Underlined langauges are not seen during the fusing process as source/target.
-A.1 Layer Analysis, Zero-shot ability and APE - AudioTower
-In our preliminary experiments, we focus exclusively on the audio modality and investigate fusing Qwen-Audio 7B (Chu et al., 2023) with Tower-Instruct 7B v0.2 (Alves et al., 2024).
-We train the fused model for ST using the same setup described earlier, but without gating, and we independently fuse the first, middle, and last layers of Qwen-Audio to assess their individual contributions.
-We then compare this fused model against several baselines.
-First, we evaluate both the out-of-the-box and fine-tuned versions of Qwen-Audio (Qwen2 Audio FT), trained end-to-end on the same ST data.
-We also include a standard cascaded setup using off-the-shelf ASR and MT models.
-Additionally, we evaluate an Automatic Post-Editing (APE) configuration in which Tower corrects translations generated by Qwen-Audio.
-To contextualize performance, we further compare against other off-the-shelf ST systems such as Tower-Spire (Ambilduke et al., 2025) and Seamless-M4T (Barrault et al., 2023), which provides insight into multilingual generalization.
-Beyond the main fused model, we also train an APE variant of our fusion architecture that post-edits the outputs of cascaded systems.
-Unlike traditional APE, this Speech-Aware APE allows the model to use the source audio when correcting errors, enabling more informed revisions.
-For both APE variants, we generate synthetic APE tuples by running Qwen-Audio on the training data and fine-tune with LoRA.
-We report results on the FLEURS test set in Table 4, including evaluations on zero-shot language directions not present on either the source or target side during training.
-First, while fine-tuning improves Qwen2 Audio, it still falls short of the out-of-the-box cascaded setup across all languages, underscoring the strength of translation-centric LLMs.
-Tower-based APE is competitive for seen languages but struggles in harder zero-shot directions, such as frkr.
-Off-the-shelf end-to-end ST models lag significantly behind cascaded systems as well.
-Our proposed fused model (AudioTower) with middle-layer fusion achieves the strongest overall performance and even surpasses the cascaded baseline in several languages.
-Notably, both the first and middle layers show strong transfer to unseen languages, whereas the last layer collapses entirely—supporting the hypothesis that upper layers encode highly language-specific features, making them less suitable for cross-lingual contextual embeddings.
-Finally, the APE variant of the fused model delivers the best results overall by combining the strong hypotheses of the cascaded pipeline with the ability to exploit the raw audio when correcting errors.
-We also test this APE model with empty hypotheses and observe performance comparable to AudioTower-Mid, indicating its flexibility for multiple use cases.
-Given the strong APE results, we additionally explore whether the model can correct outputs from different ASR systems.
-Using Whisper (Radford et al., 2023) for ASR, we evaluate the fused APE model on Whisper+Tower outputs and report results in Table˜5.
-When Whisper+Tower produce the strongest cascaded baseline, using them as hypotheses likewise yields the highest APE performance—highlighting that Speech-Aware APE can be beneficial not only for E2E setups but also for offline cascaded scenarios.
+A.1 Impact of Training Tasks
+[…]
+W/o OCR 87.12 82.42 79.75
+W/o (OCR + Self-Cascade) 89.19 84.35 82.14
+All Tasks 89.90 86.79 82.03
+All Tasks (Self-Cascade Inference) 90.05 87.20 82.47
+Table 9: Ablation study of training tasks on MCIF test set.
+To analyze the contribution of the training tasks proposed in Section 2.3, we train two ablation models: one without the OCR task and another without both OCR and self-cascading tasks.
+We evaluate all models using direct inference on the MCIF test set and report results in Table 9.
+The results reveal several insights.
+Removing the OCR task leads to the weakest performance across all language pairs, likely because the model cannot leverage the intermediate ASR step to improve direct translation.
+Removing both OCR and self-cascading tasks yields moderate performance, as this configuration cannot exploit self-cascading at inference time.
+In contrast, training on all tasks achieves the best results, particularly when combined with self-cascading inference, demonstrating that our multi-task training strategy effectively leverages both visual information and intermediate ASR outputs without causing negative interference.
 A.2 Prompt Building
@@
 Algorithm 1 Online Prompt building for examples in the training data to enable alignment and self-cascading inference mode.
+A.3 Layer Analysis, Zero-shot ability and APE - AudioTower
+In our preliminary experiments, we focus exclusively on the audio modality and investigate fusing Qwen-Audio 7B (Chu et al., 2023) with Tower-Instruct 7B v0.2 (Alves et al., 2024).
+We train the fused model for ST using the same setup described earlier, but without gating, and we independently fuse the first, middle, and last layers of Qwen-Audio to assess their individual contributions.
+We then compare this fused model against several baselines.
+First, we evaluate both the out-of-the-box and fine-tuned versions of Qwen-Audio (Qwen2 Audio FT), trained end-to-end on the same ST data.
+We also include a standard cascaded setup using off-the-shelf ASR and MT models.
+Additionally, we evaluate an Automatic Post-Editing (APE) configuration in which Tower corrects translations generated by Qwen-Audio.
+To contextualize performance, we further compare against other off-the-shelf ST systems such as Tower-Spire (Ambilduke et al., 2025) and Seamless-M4T (Barrault et al., 2023), which provides insight into multilingual generalization.
+Beyond the main fused model, we also train an APE variant of our fusion architecture that post-edits the outputs of cascaded systems.
+Unlike traditional APE, this Speech-Aware APE allows the model to use the source audio when correcting errors, enabling more informed revisions.
+For both APE variants, we generate synthetic APE tuples by running Qwen-Audio on the training data and fine-tune with LoRA.
+We report results on the FLEURS test set in Table 5, including evaluations on zero-shot language directions not present on either the source or target side during training.
+First, while fine-tuning improves Qwen2 Audio, it still falls short of the out-of-the-box cascaded setup across all languages, underscoring the strength of translation-centric LLMs.
+Tower-based APE is competitive for seen languages but struggles in harder zero-shot directions, such as frkr.
+Off-the-shelf end-to-end ST models lag significantly behind cascaded systems as well.
+Our proposed fused model (AudioTower) with middle-layer fusion achieves the strongest overall performance and even surpasses the cascaded baseline in several languages.
+Notably, both the first and middle layers show strong transfer to unseen languages, whereas the last layer collapses entirely—supporting the hypothesis that upper layers encode highly language-specific features, making them less suitable for cross-lingual contextual embeddings.
+Finally, the APE variant of the fused model delivers the best results overall by combining the strong hypotheses of the cascaded pipeline with the ability to exploit the raw audio when correcting errors.
+We also test this APE model with empty hypotheses and observe performance comparable to AudioTower-Mid, indicating its flexibility for multiple use cases.
+Given the strong APE results, we additionally explore whether the model can correct outputs from different ASR systems.
+Using Whisper (Radford et al., 2023) for ASR, we evaluate the fused APE model on Whisper+Tower outputs and report results in Table˜8.
+When Whisper+Tower produce the strongest cascaded baseline, using them as hypotheses likewise yields the highest APE performance—highlighting that Speech-Aware APE can be beneficial not only for E2E setups but also for offline cascaded scenarios.
 Category Model Name Image XCOMET-XL Minor Major Critical Total
-Cascade Systems (ASR → MT)
-Omni + SeedX
-No FT 89.66 2172 132 68 2372
+Omni FT (Direct ST) LoRA FT 85.90 2047 166 173 2386
+Cascade Systems (ASR → MT) Omni + SeedX No FT 89.66 2172 132 68 2372
 No FT 89.49 2115 133 76 2324
@@
 LoRA FT 90.18 2135 118 56 2309
-End-to-End (Direct ST)
-OmniFusion
-Mid Fusion 88.44 2126 157 74 2357
+End-to-End (Direct ST) OmniFusion Mid Fusion 88.44 2126 157 74 2357
 Mid Fusion 90.03 2256 104 68 2428
@@
 Category Model Name Image XCOMET-XL Minor Major Critical Total
-Cascade Systems (ASR → MT)
-Omni + SeedX
-No FT 86.03 850 932 51 1833
+Omni FT (Direct ST) LoRA FT 80.88 727 1018 54 1799
+Cascade Systems (ASR → MT) Omni + SeedX No FT 86.03 850 932 51 1833
 No FT 86.11 862 891 67 1820
@@
 LoRA FT 86.85 868 855 55 1778
-End-to-End (Direct ST)
-OmniFusion
-Mid Fusion 82.80 754 1065 93 1912
+End-to-End (Direct ST) OmniFusion Mid Fusion 82.80 754 1065 93 1912
 Mid Fusion 85.75 899 903 64 1866
@@
 Category Model Name Image XCOMET-XL Minor Major Critical Total
-Cascade Systems (ASR → MT)
-Omni + SeedX
-No FT 81.94 615 1255 104 1974
+Omni FT (Direct ST) LoRA FT 80.13 463 1372 44 1879
+Cascade Systems (ASR → MT) Omni + SeedX No FT 81.94 615 1255 104 1974
 No FT 82.23 597 1254 81 1932
@@
 LoRA FT 82.74 604 1326 56 1986
-End-to-End (Direct ST)
-OmniFusion
-Mid Fusion 79.47 462 1343 50 1855
+End-to-End (Direct ST) OmniFusion Mid Fusion 79.47 462 1343 50 1855
 Mid Fusion 82.15 614 1240 53 1907
@@
 Turkish tr
+Urdu ur
 Vietnamese vi
-Table 9: Language codes used in paper following Cheng et al. (2025).
+Table 13: Language codes denoted in our paper following Cheng et al. (2025).
 Hyperparameter Value
@@
 Table 15: Example illustrating where OmniFusion is better than cascaded variant. Example taken from MCIF segment 1.
+Image SeedX Output OmniFusion Output Reference
+[Uncaptioned image] AUSGANG AUSFAHRT AUSFAHRT
+[Uncaptioned image] AUSGANG AUSGANG AUSGANG
+Table 16: Example illustrating OmniFusion disambiguates source sentence with image context.
+The word ’EXIT’ when translated to German can be ’AUSFAHRT’ (as vehicle exit) or ’AUSGANG’ (as person exit) and needs the image for disambiguation.